### v2

# Estimating non-linear functionals of trawl processes

Abstract
Trawl processes is a family of continuous-time, infinitely divisible, stationary processes whose correlation structure is entirely characterized by its so-called trawl function. This paper investigates the problem of estimating non-linear functionals of a trawl function under an in-fill and a long-span sampling scheme. Specifically, building on the work of [17], we introduce non-parametric estimators for functionals of the type $\Psi_{t}(g)=\int_{0}^{t}g(a(s))\mathrm{d}s$ and $\Lambda_{t}(g)=\int_{t}^{\infty}g(a(s))\mathrm{d}s$ where $a$ represents the trawl function of interest and $g$ a non-linear test function. We show that our estimator for $\Psi_{t}(g)$ is consistent and asymptotically Gaussian regardless of the memory of the process. We further demonstrates that the same phenomenon occurs for the estimation of $\Lambda_{t}(g)$ as long as $g(x)=\mathrm{O}(\lvert x\rvert^{p})$, as $x\to 0$, for some $p>3$. Additionally, we illustrate how our results can be used to test the presence of $T$-dependent data that is robust to persistence.
1 Introduction
Trawl processes form a subclass of continuous-time, strictly stationary, and infinitely divisible processes. Specifically, a trawl process $X=(X_{t})_{t\in\mathbb{R}}$ is constructed by evaluating a Lévy basis – also known as infinitely divisible independently scattered random measure – $L$ over a set $A_{t}$, known as trawl set, which is typically a subset of $\mathbb{R}^{2}$. A trawl process, by its design, offers highly flexible autocorrelation structure and the ability to produce a wide range of marginal distributions, all within the framework of infinitely divisible distributions. In fact, the distribution of $X_{t}$ is completely determined by $L$, while the autocorrelation function of $X$ is fully described by the so-called trawl function $a$. This relationship is often expressed through the identity
$\rho_{X}(h)=\frac{\int_{h}^{\infty}a(s)\mathrm{d}s}{\int_{0}^{\infty}a(s)\mathrm{d}s},\,\,\,h\geq 0.$
These features make trawl processes suitable for modeling data with stylized facts such as non-Gaussianity, heavy tails, jumps, and persistence. This is reflected by their increasingly popularity for modeling complex temporal phenomena across disciplines such as finance ([3, 4, 20, 23, 24]) and physics ([9, 14, 18]).
Regarding statistical inference for trawl processes, various parametric approaches have been explored in the literature. See for instance [1, 3, 4, 19]. In this context, the work in [17] represents the first to address non-parametric estimation of the trawl function and serves as the starting point for the present study. Specifically, the authors in that paper showed that, given $n$ equidistant observations of $X$, say $(X_{i\Delta_{n}})_{i=0}^{n}$, the trawl function $a$ can be consistently estimated by
$\check{a}(t)$ $=-\Delta_{n}^{-1}\left[\hat{\varGamma}_{l+1}-\hat{\varGamma}_{l}\right],\,\,\,\text{if }\Delta_{n}l\leq t<(l+1)\Delta_{n},$ (1)
assuming that as $n\to\infty$, $\Delta_{n}\to 0$, with $n\Delta_{n}\to\infty$, and provided the Lévy seed of $L$ (see Section 2) has unit variance. $\hat{\varGamma}_{l}$ in (1) denotes the sample auto-covariance function of $X$ at lag $l\Delta_{n}$. Furthermore, under suitable conditions, the estimator $\check{a}(t)$ is asymptotically normal with asymptotic variance given by
$\sigma_{a}^{2}(t):=$ $\mathfrak{K}_{4}a(t)+2\int_{0}^{\infty}a(s)^{2}\mathrm{d}s+2\int_{0}^{t}a(t-s)a(t+s)\mathrm{d}s$
$-2\int_{t}^{\infty}a(s-t)a(t+s)\mathrm{d}s,$
for some non-negative constant $\mathfrak{K}_{4}$. In the same work, the authors proposed to use $\check{a}$ once again along with a “Riemann sum” approach to estimate the quadratic functionals of $a$ appearing in the definition of $\sigma_{a}^{2}(t)$, e.g. $\int_{0}^{\infty}a(s)^{2}\mathrm{d}s$. This procedure has some limitations. First, it relies on a tuning parameter whose choice and finite-sample performance are unclear. Second, the estimation error of $\int_{0}^{\infty}a(s)^{2}\mathrm{d}s$ is dependent of the behaviour of $\check{a}(t)$ for $t$ large, which could result on a statistic that is very sensitive to the memory of the underlying stochastic process. In this paper, we address these issues by considering the general problem of estimating functionals of the type
$\Psi_{t}(g)=\int_{0}^{t}g(a(s))\mathrm{d}s;\,\,\Lambda_{t}(g):=\int_{t}^{\infty}g(a(s))\mathrm{d}s,$
where $g$ is a continuous non-linear test function, based on the sample $(X_{i\Delta_{n}})_{i=0}^{n}$.
Main contributions
By slightly modifying the definition of $\check{a}$ in (1), which we denote by $\hat{a}$ (see (9) below), we derive limit theorems for the functionals
$\Psi_{t}^{n}(g):=$ $\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}g(\hat{a}(l\Delta_{n})),\,\,\,\,0\leq t\leq(n-1)\Delta_{n};$
$\Lambda_{t}^{n}(g):=$ $\Delta_{n}\sum_{l=[t/\Delta_{n}]}^{n-1}g(\hat{a}(l\Delta_{n}));\,\,0\leq t\leq(n-1)\Delta_{n},$
using a double-asymptotic scheme based on both long-span and high-frequency observations of the underlying trawl process $X$. The function $g$ is assumed to be continuous and of polynomial growth.
Under minimal integrability conditions on $X$, we provide a functional law of large numbers for $\Psi^{n}(g)$. With stronger assumptions on the trawl function, we establish a functional Central Limit Theorem (CLT from now on) at the rate $\sqrt{n\Delta_{n}}$, applicable to test functions that are twice continuously differentiable. Additionally, in line with [17], we show that this convergence is independent of the memory of $X$. In contrast, the asymptotic behaviour of $\Lambda^{n}(g)$ is more subtle. First, we demonstrate that, in general, $\Lambda_{t}^{n}(g)$ is not consistent for $\int_{t}^{\infty}g(a(s))\mathrm{d}s$ unless $g(x)=\mathrm{O}(\lvert x\rvert^{p})$, as $x\to 0$, for some $p>2$. In the latter situation with $p>3$, $\Lambda_{t}^{n}(g)$ also fulfils a CLT analogous to that of $\Psi^{n}(g)$. Secondly, to mitigate the inconsistency of $\Lambda^{n}(g)$ we alter its definition (see equation (13)) following the approach suggested in [17]. The resulting estimator $\bar{\Lambda}^{n}(g)$ is consistent and satisfies a CLT at rate $\sqrt{n\Delta_{n}}$. However, this modification requires a tunning parameter and its second-order asymptotics are now dependent on the memory of the process and the sample scheme. Finally, to illustrate the usefulness of our results, we propose a method for testing the presence of $T$-dependent data that is robust to long memory.
Outline of the paper
The remainder of the paper is organized as follows. In Section 2 we introduce the main mathematical concepts and and notation that will be used through the paper. Furthermore, the definitions of trawl processes and its basic properties are also included in this part. In Section 3 we formulate the main results of the paper, concerning the asymptotics of the functionals $\Psi^{n}(g)$ and $\Lambda^{n}(g)$, along with our new test for $T$-dependence. Due to the technical nature of our proofs, most of them are deferred to Section 4.
2 Preliminaries
This part focus on introducing the basic notation and concepts that will be used during this paper. Throughout the following sections $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ will denote a complete probability space. As usual, a function $f:\mathbb{R}\rightarrow\mathbb{R}$ is said to be of class $\mathcal{C}^{N}$, $N\in\mathbb{N}$, if it is $N$-times continuously differentiable. In this case, $f^{(k)}$ will denote the $k$th derivative of $f$ with the convention that $f^{(0)}=f$. Often, we will write $f^{\prime}$ or $\dot{f}$ instead of $f^{(1)}$. The symetrization of a function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$ is defined and denoted as $\tilde{f}(x,y):=f(x,y)+f(y,x)$.
[…]
We recall to the reader that a real-valued random field $L=\{L\left(A\right):A\in\mathcal{B}_{b}(\mathbb{R}^{d})\}$ on $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ is called a homogeneous Lévy basis if it is an infinitely divisible (ID for short) independently scatterred random measure such that
$\mathbb{E}(\exp(\mathbf{i}zL(A)))=\exp(\mathrm{Leb}(A)\psi(z)),\,\,\forall\,A\in\mathcal{B}_{b}(\mathbb{R}^{d}),z\in\mathbb{R},$ (2)
where $\mathrm{Leb}$ and $\mathcal{B}_{b}(\mathbb{R}^{d})$ denote the Lebesgue measure on $\mathbb{R}^{d}$ and the bounded Borel sets of $\mathbb{R}^{d}$, respectively. Furthermore,
[…]
with $\gamma\in\mathbb{R},$ $b\geq 0$ and $\nu$ a Lévy measure, i.e. $\{$$\int_{\mathbb{R}\backslash\{0\}}(1\land\rvert x\rvert^{2})\nu(\mathrm{d}x)<\infty$ and $\nu$ charges 0 at 0{0}. We will refer to the ID random variable with characteristic triplet $\left(\gamma,b,\nu\right)$ as the Lévy seed of $L$, and it will be denoted by $L^{\prime}$. As it is costumary, $\left(\gamma,b,\nu\right)$ will be called the characteristic triplet of $L$ and $\psi$ its characteristic exponent. We will say that $L$ has unit variance if $\mathrm{Var}(L^{\prime})=1$.
Let $L$ be a homogeneous Lévy basis on $\mathbb{R}^{2}$ with characteristic triplet $(\gamma,b,\nu)$. In addition, we set
$A=\left\{(r,y):r\leq 0,0\leq y\leq a(-r)\right\}.$ (4)
where $a:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$ is a non-increasing, continuous, and integrable function. The process following the dynamics
$X_{t}:=L(A_{t}),\,\,\,t\in\mathbb{R},$ (5)
where $A_{t}:=A+(t,0)$, is called trawl process. To avoid trivial situations, we will always assume that $\mathrm{Leb}(A)>0$ or equivalently $a$ cannot be equal to 0 almost everywhere. From now on, we will allude to $A$ and $a$ as trawl set and trawl function, respectively. Furthermore, $L$ will termed as as the background driving Lévy basis.
It is well known that $X$ is strictly stationary and, in the case when $L^{\prime}$ is square integrable, its auto-covariance function is given by
$\varGamma_{X}(h):=\mathrm{Var}(L^{\prime})\int_{\left|h\right|}^{\infty}a(u)\mathrm{d}u,\,\,\,h\in\mathbb{R}.$ (6)
For a detailed exposition on the basic properties of trawl processes and Lévy bases we refer the reader to [2] and references therein.
3 Estimating non-linear functionals of trawl functions
In what is left of this work $X$ will denote a trawl process with trawl function $a$ and whose background driving Lévy basis has unit variance. As already discussed in the introduction, our main goal is to estimate the functionals
$\Psi_{t}(g)=\int_{0}^{t}g(a(s))\mathrm{d}s;\,\,\Lambda_{t}(g):=\int_{t}^{\infty}g(a(s))\mathrm{d}s,$ (7)
when equidistant observations of $X$, say $(X_{i\Delta_{n}})_{i=0}^{n}$, are available. The sample scheme considered in this work is in-fill and long-span, i.e.
$n\uparrow\infty;\,\,\Delta_{n}\downarrow 0;\,\,n\Delta_{n}\rightarrow+\infty.$ (8)
To do so, we modify slightly the definition of $\check{a}$ in (1) as follows:
$\hat{a}(t):=-\frac{1}{n\Delta_{n}}\sum_{k=l}^{n-1}(X_{(k-l)\Delta_{n}}-\bar{X}_{n})\delta_{k}X,\,\,l\Delta_{n}\leq t<(l+1)\Delta_{n},$ (9)
for some $l=0,\ldots,n-1$ and where $\delta_{k}X:=X_{(k+1)\Delta_{n}}-X_{k\Delta_{n}}$. It is not difficult to see that $\hat{a}(t)$ and $\check{a}(t)$ are asymptotically equivalents. The estimators for the functionals appearing in (7) are constructed in a natural way as
$\Psi_{t}^{n}(g)$= $\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}g(\hat{a}(l\Delta_{n})),0\leq t\leq(n-1)\Delta_{n};$ (10)
$\Lambda_{t}^{n}(g)$= $\Delta_{n}\sum_{l=[t/\Delta_{n}]}^{n-1}g(\hat{a}(l\Delta_{n}));0\leq t\leq(n-1)\Delta_{n}.$
3.1 Limit theorems for $\Psi^{n}(g)$ and $\Lambda^{n}(g)$
In this part we describe first and second order asymptotics for the estimators introduced in (10). We start discussing consistency of such statistics. We recall the reader that a real-valued function $f:\mathbb{R}\rightarrow\mathbb{R}$ is said to be of polynomial growth of order $q\geq 0$ if there is some $C>0$, such that
$\rvert f(x)\rvert\leq C(1+\rvert x\rvert^{q}),\,\,\forall\,x\in\mathbb{R}.$
The reader should keep in mind that our sampling scheme satisfies (8) and this will be always assumed to hold for the rest of this paper. Under this set-up we have the following consistency result for $\Psi^{n}(g)$.
Theorem 1.
Let $g:\mathbb{R}\rightarrow\mathbb{R}$ be a continuous function of polynomial growth of order $q\geq 0$. Suppose that $\mathbb{E}(\rvert L^{\prime}\rvert^{2q\lor 4})<\infty$. Then, $\Psi^{n}(g)\overset{u.c.p}{\rightarrow}\Psi(g).$
Let us now turn our attention to $\Lambda^{n}(g)$. To guarantee integrability of the mapping $s\mapsto g(a(s))$, ergo the well-definiteness of $\Lambda_{t}(g)$, we need to restrict the type of test functions to be considered. For a given $p,q\geq 0$, $d\in\mathbb{N}$, we will write $\mathfrak{C}_{p,q}^{d}$ for the family of functions of class $\mathcal{C}^{d}$ satisfying that $g^{(j)}(0)=0$ for $j=0,1,\ldots,d-1$ and that $g^{(d)}(x)=\mathrm{O}(\rvert x\rvert^{p})$ as $x\rightarrow 0$ and $g^{(d)}(x)=\mathrm{O}(\rvert x\rvert^{q})$ as $x\rightarrow+\infty$. Obviously the power function $g(x)=\rvert x\rvert^{\alpha}$, for $\alpha>1$ belongs to $\mathfrak{C}_{p,q}^{[\alpha]}$ with $p=q=\alpha-[\alpha]$.
Remark 1.
Note that if $g\in\mathfrak{C}_{p,q}^{d}$, an induction argument along with the Mean-Value Theorem imply that $g^{(j)}(x)=\mathrm{O}(\rvert x\rvert^{d-j+p})$, as $x\rightarrow 0$, for all $j=1,\ldots,d$. Thus, in view that $a$ is decreasing and bounded by $a(0),$ we can find a constant $C=C(a(0))>0$, such that for all $s\geq 0$
$\max_{j=0,1,\ldots,d-1}\rvert g^{(j)}(a(s))\rvert\leq Ca(s)^{d-j+p},\,\,\,j=1,\ldots,d.$ (11)
It follows that the mappings $s\in[0,+\infty)\mapsto g^{(j)}(a(s))$ are continuous, bounded, and integrable for $j=0,1,\ldots,d\land(d-1+[p])$. In particular, $\Lambda_{t}(g)$ is well-defined for every $t\geq 0$.
Theorem 2.
Let $g\in\mathfrak{C}_{p,q}^{1}$ with $p,q>1$ and assume that $\mathbb{E}(\rvert L^{\prime}\rvert^{2(q\lor p+1)})<\infty$. Then, as $n\rightarrow\infty$, $\Lambda^{n}(g)\overset{u.c.p}{\rightarrow}\Lambda(g).$
The previous result does not cover the quadratic case, i.e. when $p\land q=1$. This is because in this situation $\Lambda^{n}(g)$ might be biased as the following result shows.
Theorem 3.
Let $g(x)=x^{2}$ and assume that $\mathbb{E}(\rvert L^{\prime}\rvert^{4})<\infty$. Then, as $n\rightarrow\infty$, $\Lambda_{0}^{n}(g)\overset{\mathbb{P}}{\rightarrow}2\Lambda_{0}(g).$
To deal with the previous situation, we proceed as in [17] and modify $\Lambda^{n}(g)$ as follows: We introduce a sequence $N_{n}\in\mathbb{N}$, $N_{n}\leq n$, such that
$i)\,N_{n}\uparrow\infty;\,\,ii)\,N_{n}/n\rightarrow 0;\,\,iii)\,N_{n}\Delta_{n}\rightarrow+\infty.$ (12)
We set
$\bar{\Lambda}_{t}^{n}(g):=\Delta_{n}\sum_{l=[t/\Delta_{n}]}^{N_{n}-1}g(\hat{a}(l\Delta_{n}));\,\,0\leq t\leq(N_{n}-1)\Delta_{n}.$ (13)
The sequence of processes $\bar{\Lambda}^{n}$, unlike $\Lambda^{n}$, is consistent in the quadratic case. More precisely:
Theorem 4.
Consider $N_{n}\in\mathbb{N}$ as in (12). Let $p,q\geq 1$ and assume that $\mathbb{E}(\rvert L^{\prime}\rvert^{2(q\lor p+1)})<\infty$. Then for all $g\in\mathfrak{C}_{p,q}^{1}$, $\bar{\Lambda}_{t}^{n}(g)\overset{u.c.p}{\rightarrow}\Lambda(g)$, as $n\rightarrow\infty.$
We now proceed to discuss second order asymptotics for our statistics. As usual, to achieve this, stronger assumptions need to be imposed. We follow essentially the set-up of [17].
Assumption 1.
The sampling scheme satisfies that $n\Delta_{n}^{3}\rightarrow 0$ as $n\rightarrow\infty$. Furthermore, the trawl function admits the representation $a(s)=\int_{s}^{\infty}\phi(y)dy$, for $s\geq 0$, where $\phi$ is a strictly positive càdlàg function such that $\phi(s)=\mathrm{O}(s^{-\alpha-1})$ as $s\rightarrow+\infty$ for some $\alpha>1$.
Before presenting the Central Limit Theorems for $\Psi^{n},\Lambda^{n}$, and $\bar{\Lambda}^{n}$ let us introduce some functions that will serve for describing their asymptotic variance. For $s,r\geq 0$ and $\mathfrak{K}_{4}:=\int x^{4}\nu(dx)$, we put
$\sigma_{1}(s,r)$ $:=\mathfrak{K}_{4}a(s\lor r);$
$\sigma_{2}(s,r)$ $:=\int_{0}^{\infty}a(u)a(\rvert u-(s-r)\rvert)\mathrm{sgn}(u-(s-r))\mathrm{d}u;$
$\sigma_{3}(s,r)$ $:=\int_{0}^{\infty}a(u+r)a(\rvert s-u\rvert)\mathrm{sgn}(s-u)\mathrm{d}u.$
Theorem 5.
Let Assumption 1 hold. If $g:\mathbb{R}\rightarrow\mathbb{R}$ is of class $\mathcal{C}^{2}$, with $g^{(2)}$ being of polynomial growth of order $q\geq 0$ and $\mathbb{E}(\rvert L^{\prime}\rvert^{2(q+2)\lor r_{0}})<\infty$ for some $r_{0}>4$, then as $n\rightarrow\infty$
$\sqrt{n\Delta_{n}}(\Psi^{n}(g)-\Psi(g))\overset{\mathcal{D}([0,T])}{\Longrightarrow}Z,$
where $(Z_{t})_{t\geq 0}$ is a centered Gaussian process with
$\mathbb{E}(Z_{t}Z_{s})=\int_{0}^{t}\int_{0}^{s}\dot{g}(a(u))\Sigma_{a}(u,r)\dot{g}(a(r))\mathrm{d}r\mathrm{d}u,\,\,\,t,s\geq 0,$
in which
$\Sigma_{a}(u,r)$ $=\sigma_{1}(u,r)+\tilde{\sigma}_{2}(u,r)+\tilde{\sigma}_{3}(u,r).$ (14)
The second order asymptotics for the functionals $\Lambda^{n}$ and $\bar{\Lambda}^{n}$ will be much alike to the one presented in the previous theorem. However, we will need to impose certain restrictions on the window $N_{n}$ that will help us to control the discretization error associated to $\bar{\Lambda}^{n}$ as well as the memory of $X$. Specifically, we will assume that:
Assumption 2.
There are $1<\varpi<3,\,\theta>0$, and $0<\kappa<1$, such that $n\Delta_{n}^{\varpi}$ is bounded from above and below, and $N_{n}\sim\theta n^{\kappa}$, as $n\rightarrow\infty$.
The CLT for the functionals $\Lambda^{n}$ and $\bar{\Lambda}^{n}$ reads as follows.
Theorem 6.
Let Assumption 1 hold. Suppose that $g\in\mathfrak{C}_{p,q}^{2}$ and that for some $r_{0}>4$ we have that $\mathbb{E}(\rvert L^{\prime}\rvert^{2(q\lor p+2)\lor r_{0}})<\infty$. Then the following holds:
  1. 1.
    If $p,q>1$, then as $n\rightarrow\infty$
    $\sqrt{n\Delta_{n}}(\Lambda^{n}(g)-\Lambda(g))\overset{\mathcal{D}([0,T])}{\Longrightarrow}Z,$ (15)
    where $(Z_{t})_{t\geq 0}$ is a centered Gaussian process with
    $\mathbb{E}(Z_{t}Z_{s})=\int_{t}^{\infty}\int_{s}^{\infty}\dot{g}(a(u))\Sigma_{a}(u,r)\dot{g}(a(r))\mathrm{d}r\mathrm{d}u,$
    with $\Sigma_{a}$ as in (14).
  2. 2.
    If in addition Assumption 2 is satisfied with
    $0<\frac{1}{\varpi}+\frac{\varpi-1}{2\varpi((2+p)\alpha-1)}<\kappa<\frac{1}{2}(1+\frac{1}{\varpi}),$ (16)
    and $p,q\geq 0$ arbitrary, then (15) also holds if we replace $\Lambda^{n}$ by $\bar{\Lambda}^{n}$.
Remark 2.
The following remarks are in order
  1. 1.
    Let $Z$ be as in Theorems 5 and 6. From Remark 1, it follows that for every $0\leq s\leq t$
    $\mathbb{E}(\rvert Z_{t}-Z_{s}\rvert^{2})\leq C(t-s)^{2}.$
    Therefore, by Kolmogorov’s criterion, we deduce that $Z$ has $\lambda$-Hölder continuous paths for every $\lambda<1.$
  2. 2.
    We are able to show that in the limiting case $p\wedge q=1$ – recall that when this is not the case, in general, $\Lambda^{n}(g)$ is not consistent – the sequence $\sqrt{n\Delta_{n}}(\Lambda^{n}(g)-\Lambda(g))$ is tight on $\mathcal{D}([0,T])$ but we were not able to identify its limit. However, we believe that, just as in the case $g(x)=x^{2}$, the limit process is Gaussian but no longer centred.
  3. 3.
    The lower bound in (16), which depends on the memory of $X$, is to prevent that the tail $\int_{N_{n}\Delta_{n}}^{\infty}g(a(s))\mathrm{d}s$ dominates the asymptotics. Thus, unlike Theorem 5 and the first part of Theorem 6, the CLT for $\bar{\Lambda}^{n}(g)$ is no longer robust to long memory.
3.2 A test for $T$-dependent trawls
In this part we briefly exemplify how our results can be used to test for $T$-dependence in a non-parametrical framework. We recall the reader that a stochastic process $(Y_{t})_{t\in\mathbb{R}}$ is said to be $T$-dependent, with $T>0$, if for a given $t_{0}\in\mathbb{R}$, the processes $(Y_{t})_{t\leq t_{0}}$ and $(Y_{t})_{t\geq T+t_{0}}$ are independent (in the sense of finite-dimensional distributions). We have the following characterization of $T$-dependence for trawl processes.
Proposition 1.
Let $X$ be a trawl process. Then, unless $L$ is deterministic, $X$ is $T$-dependent if, and only if, for all $t\geq T$, $\mathrm{Leb}(A_{t}\cap A)=0$.
Proof.
Since $X$ is an infinitely divisible process, $T$-dependence is equivalent to having that for a given $t_{0}$, $X_{s}$ and $X_{u}$ are independent for all $s\leq t_{0}$ and $u\geq t_{0}+T$ (see for instance Exercise E 12.9 in [16]). Furthermore, by stationarity, the former condition is equivalent to the independence of $X_{t}$ and $X_{0}$ for all $t\geq T$. Now note that, in view of (2) and the independent scatterred property of $L$, the characteristic exponent of the vector $(X_{0},X_{t})$ equals to
$\psi(z_{1})+\psi(z_{2})]\mathrm{Leb}(A)+\mathrm{Leb}(A_{t}\cap A)[\psi(z_{1}+z_{2})-\psi(z_{1})-\psi(z_{2})],\,\,\forall\,(z_{1},z_{2})\in\mathbb{R}^{2}.$
In consequence, $X_{t}$ and $X_{0}$ are independent if, and only if,
$\mathrm{Leb}(A_{t}\cap A)[\psi(z_{1}+z_{2})-\psi(z_{1})-\psi(z_{2})]=0,\,\,\forall\,(z_{1},z_{2})\in\mathbb{R}^{2}.$ (17)
Thus, if $\mathrm{Leb}(A_{t}\cap A)=0$, for all $t\geq T$, the previous relation is trivially fulfilled. Reciprocally, if (17) is satisfied for all $t\geq T$, then either $\mathrm{Leb}(A_{t}\cap A)=0$ or $\psi(z_{1}+z_{2})=\psi(z_{1})+\psi(z_{2})$ for all $(z_{1},z_{2})\in\mathbb{R}^{2}$. If the latter holds, we would have that $\psi$ is continuous and additive, so linear (c.f. Cauchy’s functional equation) which is the case only when $L$ is deterministic. ∎
The preceeding result shows that, under our set-up, $X$ is $T$-dependent if, and only if, $\int_{T}^{\infty}a(s)\mathrm{d}s=0$. Note that this condition is in turn equivalent to $\mathrm{Cov}(X_{T},X_{0})=0$ and thus one can use the sample autocovariance function to test $T$-dependence on $X$. However, it is well known that such a statistic is heavily dependent on the memory of the process. See for instance [6, 7, 10, 21]. To avoid this problem we propose to use another characterization of $T$-dependence, namely
$\int_{T}^{\infty}\rvert a(s)\lvert^{p}\mathrm{d}s=0,\text{ for some (and thus for all) }p>0,$ (18)
which follows from the fact that $a$ is non-negative and non-increasing. Therefore, under the null hypothesis of $T$-dependence we deduce, due to the fact that $\mathrm{Leb}(A)>0$, that
$\frac{\Lambda_{0}(g)}{\Psi_{T}(g)}=\frac{\int_{0}^{\infty}\rvert a(s)\lvert^{p}\mathrm{d}s}{\int_{0}^{T}\rvert a(s)\lvert^{p}\mathrm{d}s}=1,$
where $g(x)=\lvert x\rvert^{p}$. Based on the previous discussion, we proposed to use the “memory independent” statistic
$\tau_{n,p}=\frac{\Lambda^{n}_{0}(g)}{\Psi_{T}^{n}(g)}-1=\frac{\Delta_{n}\sum_{l=0}^{n-1}\rvert\hat{a}(l\Delta_{n})\rvert^{p}}{\Delta_{n}\sum_{l=0}^{[T/\Delta_{n}]-1}\rvert\hat{a}(l\Delta_{n})\rvert^{p}}-1,\,\,p>3,$
to test the hypothesis (18). The asymptotic behaviour of $\tau_{n,p}$ is described in the following result.
Corollary 1.
Let Assumption 1 hold and assume that $\mathbb{E}(\rvert L^{\prime}\rvert^{2p})<\infty$ for some $p>3$. Then,
[…]
Proof.
Plainly, $\tau_{n,p}=\Lambda^{n}_{T}(g)/\Psi_{T}^{n}(g)$, with $g(x)=\lvert x\rvert^{p}$. Thus, if (18) is satisfied, Theorem 6 implies that $\sqrt{n\Delta_{n}}\Lambda^{n}_{T}(g)\overset{d}{\to}\zeta$, where $\zeta$ is a centred Gaussian random variable with variance
$p^{2}\int_{T}^{\infty}\int_{T}^{\infty}\rvert a(u)\rvert^{p-1}\Sigma_{a}(u,r)\rvert a(r)\rvert^{p-1}\mathrm{d}u\mathrm{d}r=0,$
thanks to (18). This shows the first part of (19). In constrast, if (18) does not hold then necessarily $\int_{0}^{T}\rvert a(s)\lvert^{p}\mathrm{d}s>0$. Theorem 2 now gives that $\tau_{n,p}\overset{\mathbb{P}}{\to}\frac{\int_{T}^{\infty}\rvert a(s)\lvert^{p}\mathrm{d}s}{\int_{0}^{T}\rvert a(s)\lvert^{p}\mathrm{d}s}>0$, from which the second part of (19) follows. ∎
It should be clear from our discussion above and the previous corollary that a large value for $\sqrt{n\Delta_{n}}\tau_{n,p}$ suggests that the data is unlikely to be $T$-dependent.
Remark 3.
As pointed out in Remark 2, our proposed statistics for testing (18) is robust to the memory of $X$. Furthermore, as demonstrated in the previous corollary, $\tau_{n,p}$ is more efficient under null. However, it still suffers of some weakness. First, to rigorously use $\tau_{n,p}$ for testing $T$-dependence, a non-trival CLT under the null hipothesis (18) is required. Secondly, it remains to be seen how sensitive is $\tau_{n,p}$ in finite-sample for different values of $p$. These problems requires additional analysis, which we postpone to future research.
4 Proofs
Throughout all our proofs the non-random positive constants will be denoted by the generic symbol $C>0$, and they may change from line to line. We will often write $a\lesssim_{\theta}b$ whenever $a\leq C_{\theta}b$ where the constant $C_{\theta}$ only depends on a given parameter $\theta$. Next, for $i,j=0,1,\ldots$ and $j\geq i$ define
$\mathcal{P}_{A}^{n}(i,j):=\{(r,s):a(t_{j+1}-s)<r\leq a(t_{j}-s),t_{i-1}<s\leq t_{i}\},$ (20)
where $t_{i}=i\Delta_{n}$ with the convention that $t_{-1}=$-$\infty$. Note that for all $k=0,1,\ldots,$
$\mathrm{Leb}(A_{t_{k}}\backslash\bigcup_{i=0}^{k}\bigcup_{j\geq k}\mathcal{P}_{A}^{n}(i,j))=0.$ (21)
Using these sets we introduce the filtrations:
$\mathscr{F}_{k}^{n}:=\sigma\left(\{L(\mathcal{P}_{A}^{n}(m,j))\}_{j\geq m}:m=0,\ldots,k\right),\,\,\,\,k\geq 0,$ (22)
and
$\mathscr{G}_{j}^{n}:=\sigma\left(\{L(\mathcal{P}_{A}^{n}(i,m))\}_{0\leq i\leq m}:m=0,\ldots,j\right),\,\,\,j\geq 0.$ (23)
In what is left of this section we set $\mathfrak{t}_{L}=(\gamma,\sigma,\nu)$, the characteristic triplet of $L$, and use the notation $\tau_{n}(s)=[s/\Delta_{n}]\Delta_{n}$ as well as $u_{n}(s)=\left\lceil s/\Delta_{n}\right\rceil\Delta_{n}$, for $s\geq 0.$
4.1 Preliminary estimates and decompositions
In this part we derive some estimates and establish some fundamental decompositions that will be used in a later stage. We also introduce several random variables and processes.
The following estimate for the moments of a homogeneous Lévy basis will be fundamental in our analysis. For a proof see [13] as well as [22] (Corollary 1.2.7).
Lemma 1.
Let $L$ be a centered homogeneous Lévy basis with characteristic triplet $\mathfrak{t}_{L}=(\gamma,\sigma,\nu)$ with $\mathbb{E}(\rvert L^{\prime}\rvert^{r})<\infty$ for $r\geq 1$. Then
$\mathbb{E}(\rvert L(B)\rvert^{r})\leq C_{r,\mathfrak{t}_{L}}\mathrm{Leb}(B)^{r/2}\lor\mathrm{Leb}(B),\,\,\forall\,B\in\mathcal{B}(\mathbb{R}^{2}),$ (24)
where $C_{q,\mathfrak{t}_{L}}>0$ only depends on $q$ and $\mathfrak{t}_{L}$. Furthermore
$\mathbb{E}(L(A)^{2})$ $=\mathrm{Leb}(A)(\sigma^{2}+\int x^{2}\nu(dx)).$
$\mathbb{E}(L(A)^{3})$ $=\mathrm{Leb}(A)\int x^{3}\nu(dx).$
$\mathbb{E}(L(A)^{4})$ $=\mathrm{Leb}(A)\left\{\mathfrak{K}_{4}(L^{\prime})+3\mathrm{Leb}(A)(\sigma^{2}+\int x^{2}\nu(dx))^{2}\right\}.$
Moreover, if $A\subseteq B\subseteq C\subseteq D$, then
$\mathbb{E}\left[L(A)L(B)L(C)\right]=\mathbb{E}\left[L(A)^{3}\right],$
and
[…]
$+2\mathbb{E}[L(A)^{2}]\mathbb{E}[L(B\backslash A)^{2}]$
Next, we observe that the estimation error due to $\hat{a}$ can be represented as
$\hat{a}(l\Delta_{n})-a(l\Delta_{n})$ $=\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))+\mathbb{E}(\tilde{a}(l\Delta_{n}))-a(l\Delta_{n})$ (25)
$+\frac{1}{n\Delta_{n}}(X_{n\Delta_{n}}-X_{l\Delta_{n}})\frac{1}{n}S_{n}$
$=$ $\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))+(X_{n\Delta_{n}}-X_{l\Delta_{n}})\frac{1}{n^{2}\Delta_{n}}S_{n}$
$+\frac{(n-l)}{n}\frac{1}{\Delta_{n}}\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}[a(s)-a(l\Delta_{n})]\mathrm{d}s-\frac{l}{n}a(l\Delta_{n}),$
where, $S_{n}=\sum_{k=0}^{n-1}\tilde{X}_{k\Delta_{n}}$ and
$\tilde{a}(l\Delta_{n}):=-\frac{1}{n\Delta_{n}}\sum_{k=l}^{n-1}\tilde{X}_{(k-l)\Delta_{n}}\delta_{k}\tilde{X},\,\,\,l=0,\ldots,n-1,$ (26)
in which $\tilde{X}_{t}=X_{t}-\mathbb{E}(X_{t})$ and $\delta_{k}\tilde{X}=\tilde{X}_{(k+1)\Delta_{n}}-\tilde{X}_{k\Delta_{n}}$. Now, by letting $\tilde{L}(\cdot)=L(\cdot)-\mathbb{E}(L(\cdot)$, define
$\chi_{k}:=\tilde{L}(\cup_{i=0}^{k}\mathcal{P}_{A}^{n}(i,k)),$ $\,\,\,\varrho_{k}:=\tilde{L}(\cup_{j\geq k+1}\mathcal{P}_{A}^{n}(k+1,j));$ (27)
$\alpha_{k,j}:=\tilde{L}(\mathcal{P}_{A}^{n}(k,j)),\,\,\,$ $\gamma_{k}:=\tilde{L}(\cup_{j\geq n+1}\mathcal{P}_{A}^{n}(k+1,j)),$ (28)
and set
$\beta_{k,j,l}^{(1)}$ $=\tilde{L}(\cup_{i=0}^{k-l}\mathcal{P}_{A}^{n}(i,j));\,\,\,\beta_{k,j,l}^{(2)}=\tilde{L}(\cup_{i=0}^{k-l}\cup_{m=k-l}^{j}\mathcal{P}_{A}^{n}(i,m));$ (29)
$\beta_{k,l}^{(3)}$ $=\tilde{L}(\cup_{i=0}^{k-l}\cup_{m\geq n+1}\mathcal{P}_{A}^{n}(i,m)).$
Based on these random variables and (21) we have the decomposition
$\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))=$ $\frac{1}{n\Delta_{n}}\sum_{k=l}^{n-1}\left[\tilde{X}_{(k-l)\Delta_{n}}\chi_{k}-\mathbb{E}(\tilde{X}_{(k-l)\Delta_{n}}\chi_{k})\right]$ (30)
$-\frac{1}{n\Delta_{n}}\sum_{k=l}^{n-1}\tilde{X}_{(k-l)\Delta_{n}}\varrho_{k}.$
Moreover, by using that (once again due to (21))
$\tilde{X}_{(k-l_{n})\Delta_{n}}=\beta_{k,j,l}^{(2)}+\sum_{m=j+1}^{n}\beta_{k,m,l}^{(1)}+\beta_{k,l}^{(3)},$ (31)
we further have
$\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))=\frac{1}{n\Delta_{n}}\sum_{j=l}^{n}\varXi_{j,l,n}+\frac{1}{n\Delta_{n}}\sum_{j=l}^{n}R_{j,l,n},$ (32)
where $\varXi_{j,l,n}=\sum_{\ell=1}^{4}\xi_{j,l,n}^{(\ell)}$ and $R_{j,l,n}=\sum_{\ell=1}^{3}\rho_{j,l,n}^{(\ell)}$ in which
$\xi_{j,l,n}^{(1)}$ $=\left[\chi_{j}\beta_{j,j,l}^{(2)}-\mathbb{E}(\chi_{j}\beta_{j,j,l}^{(2)})\right];\,\,\,\xi_{j,l,n}^{(2)}=\sum_{k=l}^{j-1}\chi_{k}\beta_{k,j,l}^{(1)};\,\,\,\xi_{j,l,n}^{(3)}=-\sum_{k=l}^{j-1}\alpha_{k+1,j}\beta_{k,j-1,l}^{(2)};$ (33)
$\xi_{j,l,n}^{(4)}$ $=-\sum_{k=l}^{j-1}\beta_{k,j,l}^{(1)}\sum_{m=k+1}^{j}\alpha_{k+1,m}=-\sum_{k=l}^{j-1}\beta_{k,j,l}^{(1)}(\alpha_{k+1,j}+\varsigma_{k,j}),$
as well as
$\rho_{j,l,n}^{(1)}=$ $\chi_{j}\beta_{j,l}^{(3)};\,\,\rho_{j,l,n}^{(2)}=-\tilde{X}_{(j-l)\Delta_{n}}\gamma_{j};$ (34)
$\rho_{j,l,n}^{(3)}=$ $-\beta_{j,l}^{(3)}\tilde{L}(\cup_{m=j+1}^{n}\mathcal{\mathcal{P}}_{A}^{n}(j+1,m)).$
The following $r$-moments estimates were proven in [17].
Lemma 2.
Let $r\geq 2$ and suppose that $\mathbb{E}(\rvert L^{\prime}\rvert^{2r})<\infty$. Then,
$\mathbb{E}(\rvert\xi_{j,l,n}^{(1)}\rvert^{r})\lesssim_{r,a,\mathfrak{t}_{L}}\Delta_{n};\,\,\,\,\mathbb{E}(\rvert\xi_{j,l,n}^{(2)}\rvert^{r})\lesssim_{r,a,\mathfrak{t}_{L}}\Delta_{n}+\Delta_{n}\int_{0}^{n\Delta_{n}}a(s)s^{r/2-1}\mathrm{d}s,$ (35)
and
$\mathbb{E}(\rvert\xi_{j,l,n}^{(3)}\rvert^{r})\lesssim_{r,a,\mathfrak{t}_{L}}\Delta_{n};\,\,\,\,\mathbb{E}(\rvert\xi_{j,l,n}^{(4)}\rvert^{r})\lesssim_{r,a,\mathfrak{t}_{L}}\Delta_{n}+\Delta_{n}\int_{0}^{n\Delta_{n}}a(s)s^{r/2-1}\mathrm{d}s.$ (36)
With the previous estimates at hand we obtain the subsequent bounds.
Lemma 3.
Let $\tilde{a}$ be as in (26) and assume that $\mathbb{E}(\rvert L^{\prime}\rvert^{2r})<\infty$ for some $r\geq 2$. Then,
$\mathbb{E}(\rvert\tilde{a}(l\Delta_{n})-\mathbb{E}[\tilde{a}(l\Delta_{n})]\rvert^{r})\leq C_{r,a,\mathfrak{t}_{L}}/(n\Delta_{n})^{r/2},$
where $C_{r,a,\mathfrak{t}_{L}}>0$ is a constant only depending on $r$, $a$, and $\mathfrak{t}_{L}$.
Proof.
Without loss of generality we may and do assume that $L$ has mean 0 and so $X$. We will verify that
$E_{n}(l)=(n\Delta_{n})^{r/2}\mathbb{E}(\rvert\tilde{a}(l\Delta_{n})-\mathbb{E}[\tilde{a}(l\Delta_{n})]\rvert^{r}),\,\,l=0,\ldots,n-1,$
is bounded by a constant that does not depend neither on $n$ nor on $l$. By (30) and (31) we get that
$E_{n}(l)\leq 4^{r-1}\sum_{\ell=1}^{4}E_{n}^{\ell}(l),$
where
$E_{n}^{1}(l)$ $=\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}(\rvert\sum_{j=l}^{n}\xi_{j,l,n}^{(1)}\rvert^{r});\,\,\,E_{n}^{2}(l)=\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}(\rvert\sum_{j=l}^{n}\xi_{j,l,n}^{(2)}\rvert^{r});$
$E_{n}^{3}(l)$ $=\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}(\rvert\sum_{j=l}^{n}\rho_{j,l,n}^{(1)}\rvert^{r});\,\,E_{n}^{4}(l)=\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}(\rvert\sum_{j=l}^{n}X_{(j-l)\Delta_{n}}\varrho_{j}\rvert^{r}).$
Note that for every $j\geq l\geq 0$, $(\xi_{j,l,n}^{(1)},\xi_{j,l,n}^{(2)})$, $\rho_{j,l,n}^{(1)}$, and $X_{(j-l)\Delta_{n}}\varrho_{j}$ are martingale difference with respect to $\mathscr{G}_{j}^{n}$, $\mathscr{G}_{j}^{n}\lor\sigma(\beta_{j^{\prime},l^{\prime}}^{(3)})_{0\leq l^{\prime}\leq j^{\prime}\leq j},$ and $\mathscr{F}_{j}^{n}$, respectively. Thus, by Rosenthal’s inequality and Jensen’s inequality we deduce that
$E_{n}^{4}(l)$ $\lesssim_{r}\frac{1}{(n\Delta_{n})^{r/2}}\sum_{j=l}^{n}\mathbb{E}(\rvert X_{(j-l)\Delta_{n}}\varrho_{j}\rvert^{r})+\mathbb{E}\left[\left|\frac{1}{n\Delta_{n}}\sum_{j=l}^{n}(X_{(j-l)\Delta_{n}})^{2}\mathbb{E}(\varrho_{j}^{2})\right|^{r/2}\right]$ (37)
$\lesssim_{a}\frac{\mathbb{E}(\rvert X_{0}\rvert^{r})}{(n\Delta_{n})^{r/2}}\sum_{j=l}^{n}\mathbb{E}(\rvert\varrho_{j}\rvert^{r})+\mathbb{E}\left[\frac{1}{n}\sum_{j=l}^{n}\rvert X_{(j-l)\Delta_{n}}\rvert^{r}\right]$
$\lesssim_{r,a,\mathfrak{t}_{L}}\frac{\mathbb{E}(\rvert X_{0}\rvert^{r})}{(n\Delta_{n})^{r/2-1}}+\mathbb{E}(\rvert X_{0}\rvert^{r}),$
where in the last inequality we used (24). Similarly we obtain that
$\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}(\rvert\sum_{j=l}^{n}\rho_{j,l,n}^{(1)}\rvert^{r})$ $\lesssim_{r}\frac{1}{(n\Delta_{n})^{r/2}}\sum_{j=l}^{n}\mathbb{E}(\rvert\rho_{j,l,n}^{(1)}\rvert^{r})+\mathbb{E}\left[\left|\frac{1}{n}\sum_{j=l}^{n}(\beta_{j,l}^{(3)})^{2}\right|^{r/2}\right]$
$\lesssim_{r,a,\mathfrak{t}_{L}}\frac{1}{(n\Delta_{n})^{r/2-1}}+\frac{1}{n}\sum_{j=l}^{n}\mathbb{E}(\rvert\beta_{j,l}^{(3)}\rvert^{r})$
$\lesssim_{r,a,\mathfrak{t}_{L}}\frac{1}{(n\Delta_{n})^{r/2-1}}+1,$
where in the last step we also applied (21). As above we have (once again by Rosenthal’s inequality) for $i=1,2$ that
$E_{n}^{i}(l)\lesssim_{r}\frac{1}{(n\Delta_{n})^{r/2}}\sum_{j=l}^{n}\mathbb{E}(\rvert\xi_{j,n}^{(i)}\rvert^{r})+\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}\left[\left|\sum_{j=l}^{n}\mathbb{E}((\xi_{j,n}^{(i)})^{2}\mid\mathscr{G}_{j-1}^{n})\right|^{r/2}\right].$
Lemma 2 gives us that $\mathbb{E}(\rvert\xi_{j,n}^{(1)}\rvert^{r})\lesssim_{r,a,\mathfrak{t}_{L}}\Delta_{n}$ and that
$\mathbb{E}(\rvert\xi_{j,n}^{(2)}\rvert^{r})$ $\lesssim_{r,a,\mathfrak{t}_{L}}\Delta_{n}+\frac{(\Delta_{n}n)^{r/2}}{n}.$
Thus, we are left to show that for $i=1,2$
$E_{n}^{i}(l)^{\prime}=\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}\left[\left|\sum_{j=l}^{n}\mathbb{E}((\xi_{j,n}^{(i)})^{2}\mid\mathscr{G}_{j-1}^{n})\right|^{r/2}\right],$
is bounded uniformly on $n$ and $l$. From Lemma 3 in [17] and (24) we get that
$\sum_{j=l}^{n}\mathbb{E}((\xi_{j,n}^{(1)})^{2}\mid\mathscr{G}_{j-1}^{n})$ $\lesssim_{a,\mathfrak{t}_{L},r}n\Delta_{n}+\Delta_{n}\sum_{j=l+1}^{n}\left\{(\beta_{j,j-1,l}^{(2)})^{2}-\mathbb{E}[(\beta_{j,j-1,l}^{(2)})^{2}]\right\};$
$\sum_{j=l}^{n}\mathbb{E}((\xi_{j,n}^{(2)})^{2}\mid\mathscr{G}_{j-1}^{n})$ $\lesssim_{a,\mathfrak{t}_{L},r}n\Delta_{n}+\sum_{k=l+1}^{n}\chi_{k}\nu_{k}^{(1)}$
$+\sum_{k=l}^{n}[\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})]\mathrm{Leb}(\cup_{i=0}^{k-l}\cup_{j=k+1}^{n}\mathcal{P}_{A}^{n}(i,j)),$
where $\nu_{k}^{(1)}:=\sum_{k^{\prime}=l}^{k-1}\chi_{k^{\prime}}\mathrm{Leb}(\cup_{i=0}^{k^{\prime}-l}\cup_{j=k+1}^{n-1}\mathcal{P}_{A}^{n}(i,j))$. Thus,
$E_{n}^{1}(l)^{\prime}$ $\lesssim_{a,\mathfrak{t}_{L},r}1+\frac{1}{n}\sum_{j=l+1}^{n}\mathbb{E}[\rvert(\beta_{j,j-1,l}^{(2)})^{2}-\mathbb{E}[(\beta_{j,j-1,l}^{(2)})^{2}]\rvert^{r/2}]$
$\lesssim_{a,\mathfrak{t}_{L},r}2,$
and
$E_{n}^{2}(l)^{\prime}\lesssim_{a,\mathfrak{t}_{L},r}$ $1+\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}\left[\left|\sum_{k=l}^{n}\chi_{k}\nu_{k}^{(1)}\right|^{r/2}\right]$
$+\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}\left[\left|\sum_{k=l}^{n}[\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})]\mathrm{Leb}(\cup_{i=0}^{k-l}\cup_{j=k+1}^{n-1}\mathcal{P}_{A}^{n}(i,j))\right|^{r/2}\right],$
once again by means of (24). If $r>4$ we can apply twice the Rosenthal’s inequality, Jensen’s inequality, and (24) to obtain that
$E_{n}^{2}(l)^{\prime}\lesssim_{a,\mathfrak{t}_{L},r}$ $1+\frac{1}{(n\Delta_{n})^{r/2}}\sum_{k=l}^{n}\mathbb{E}(\rvert\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})\rvert^{r/2})$
$+\frac{1}{(n\Delta_{n})^{r/2}}\left|\sum_{k=l}^{n}\mathbb{E}(\rvert\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})\rvert^{2})\right|^{r/4}$
$+\frac{1}{(n\Delta_{n})^{r/2}}\sum_{k=l}^{n}\mathbb{E}(\rvert\chi_{k}\nu_{k}^{(1)}\rvert^{r/2})+\frac{1}{(n\Delta_{n})^{r/4}}\mathbb{E}\left[\left|\frac{1}{n}\sum_{k=l}^{n}(\nu_{k}^{(1)})^{2}\right|^{r/4}\right]$
$\lesssim_{a,\mathfrak{t}_{L},r}$ $1+\frac{1}{(n\Delta_{n})^{r/2-1}}+\frac{1}{(n\Delta_{n})^{r/4}}+\frac{1}{(n\Delta_{n})^{r/2-2}}$
$+\frac{1}{(n\Delta_{n})^{r/4-1}}+\frac{1}{(n\Delta_{n})^{r/4}}\frac{1}{n}\sum_{k=l}^{n}\mathbb{E}(\rvert\nu_{k}^{(1)}\rvert^{r/2})$
$\lesssim_{a,\mathfrak{t}_{L},r}$ $1+\frac{1}{(n\Delta_{n})^{r/2-1}}+\frac{1}{(n\Delta_{n})^{r/4}}+\frac{1}{(n\Delta_{n})^{r/2-2}}+\frac{1}{(n\Delta_{n})^{r/4-1}}.$
Finally, if $r\leq 4$ we can instead use the inequality $\mathbb{E}(\rvert Y\rvert^{r/2})\leq\mathbb{E}(\rvert Y\rvert^{2})^{r/4}$ along with the von Bahr-Esseen inequality to deduce that $E_{n}^{2}(l)^{\prime}$ is bounded up to a constant by
$1+\frac{1}{(n\Delta_{n})^{r/2}}\left|\sum_{k=l}^{n}\mathbb{E}[(\chi_{k}\nu_{k}^{(1)})^{2}]\right|^{r/4}+\frac{1}{(n\Delta_{n})^{r/2}}\sum_{k=l}^{n}\mathbb{E}(\rvert\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})\rvert^{r/2}),$
which is in turn uniformly bounded due to (24). ∎
We conclude this part with an estimate that will provide a way to control the approximation error in our statistics.
Lemma 4.
Let $r\geq 1,v\geq 0$ and set
$A^{n}:=\Delta_{n}\sum_{l=0}^{m_{n}-1}a(l\Delta_{n})^{v}\rvert\hat{a}(l\Delta_{n})-a(l\Delta_{n})\rvert^{r},\,\,m_{n}\in\mathbb{N},m_{n}\leq n.$
Suppose that $\mathbb{E}(\rvert L^{\prime}\rvert^{2r\lor 4})<\infty$ and that $m_{n}n\rightarrow\jmath\in(0,+\infty].$ Then,
$A^{n}=\mathrm{O}_{\mathbb{P}}((n\Delta_{n})^{-r/2}\int_{0}^{m_{n}\Delta_{n}}a(s)^{v}\mathrm{d}s)+\mathrm{o}(1).$ (38)
If Assumption 1 holds, we further have that
$A^{n}=\mathrm{O}_{\mathbb{P}}((n\Delta_{n})^{-r/2}\int_{0}^{m_{n}\Delta_{n}}a(s)^{v}\mathrm{d}s)+\mathrm{O}(\Delta_{n}^{r}).$ (39)
Proof.
From (25)
$A^{n}\lesssim_{r}$ $\Delta_{n}\sum_{l=0}^{m_{n}-1}a(l\Delta_{n})^{v}\rvert\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))\rvert^{r}$ (40)
$+\rvert\frac{1}{n}S_{n}\rvert^{r}\frac{1}{(n\Delta_{n})^{r}}\Delta_{n}\sum_{l=0}^{m_{n}-1}a(l\Delta_{n})^{v}\rvert X_{n\Delta_{n}}-X_{l\Delta_{n}}\rvert^{r}$
$+\frac{1}{(n\Delta_{n})^{r}}\int_{0}^{m_{n}\Delta_{n}}\rvert\tau_{n}(s)a(\tau_{n}(s))^{1+v/r}\rvert^{r}\mathrm{d}s$
$+\Delta_{n}\sum_{l=0}^{m_{n}-1}a(l\Delta_{n})^{v}\left(\frac{1}{\Delta_{n}}\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}[a(l\Delta_{n})-a(s)]\mathrm{d}s\right)^{r}.$
By Lemma 3 we deduce that
$\Delta_{n}\sum_{l=0}^{m_{n}-1}a(l\Delta_{n})^{v}\mathbb{E}(\rvert\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))\rvert^{r})\lesssim\frac{1}{(n\Delta_{n})^{r/2}}\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))^{v}\mathrm{d}s,$
and by the stationarity of $X$ we also obtain that
$\frac{1}{(n\Delta_{n})^{r}}\Delta_{n}\sum_{l=0}^{m_{n}-1}a(l\Delta_{n})^{v}\mathbb{E}(\rvert X_{n\Delta_{n}}-X_{l\Delta_{n}}\rvert^{r})\lesssim\frac{1}{(n\Delta_{n})^{r}}\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))^{v}\mathrm{d}s.$
Furthermore, by Jensen’s inequality and the monoticity of $a$, we deduce that the last summand in (40) is bounded by
$\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))^{v}[a(\tau_{n}(s))-a(s)]^{r}\mathrm{d}s.$ $\leq a(0)^{v}\int_{0}^{\Delta_{n}}[a(0)-a(s)]\mathrm{d}s$ (41)
$+\int_{0}^{\Delta_{n}}a(s)^{v+1}\mathrm{d}s$
$-\int_{(m_{n}-1)\Delta_{n}}^{m_{n}\Delta_{n}}a(s)^{v+1}\mathrm{d}s$
$\lesssim_{a}\Delta_{n}.$
Next, using that $a$ is integrable and of bounded variation, we can argue as in the proof of Lemma 5.4 in [15] to verify that $xa(x)\rightarrow 0$ as $x\rightarrow\infty$. Consequently
$\frac{1}{(n\Delta_{n})^{r}}\int_{0}^{m_{n}\Delta_{n}}\rvert\tau_{n}(s)a(\tau_{n}(s))^{1+v/r}\rvert^{r}\mathrm{d}s\lesssim_{r,a}\frac{1}{(n\Delta_{n})^{r}}\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))^{v}\mathrm{d}s.$
Relation (38) now follows by observing that
$\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))^{\upsilon}\mathrm{d}s\leq a(0)^{\upsilon}\Delta_{n}+\int_{0}^{m_{n}\Delta_{n}}a(s)^{v}\mathrm{d}s,$
thanks to the monoticity of $a$. Finally, if Assumption 1 is satisfied then there is $y_{0}>0$ such that
$\frac{1}{\Delta_{n}}\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}[a(l\Delta_{n})-a(s)]\mathrm{d}s$ $\leq\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}\phi(y)\mathrm{d}y$ (42)
$\leq\mathbf{1}_{l\Delta_{n}\leq y_{0}}\Delta_{n}+\mathbf{1}_{l\Delta_{n}>y_{0}}\Delta_{n}(l\Delta_{n})^{-\alpha}.$
Thus, the estimate in (41) can be replaced by
$\Delta_{n}^{r}(\mathrm{O}(1)+\int_{y_{0}}^{\infty}s^{-r\alpha}\mathrm{d}s)\leq C\Delta_{n}^{r},$
which gives (39).∎
4.2 Proof of Theorems 1 - 4
We start by showing the consistency of $\Psi^{n}(g)$. The proof follows the idea of the proof of Theorem 9.4.1 in [11] along with Lemma 3. For the rest of this section we fix $T>0$ and choose $n$ large enough such that $(n-1)\Delta_{n}>T.$
Proof of Theorem 1.
Without loss of generality we may and do assume that $g\geq 0$ (otherwise analyze $g^{+}$ and $g^{-}$ separately). Thus, we only need to check that for every $0\leq t\leq T$, $\Psi_{t}^{n}(g)\overset{\mathbb{P}}{\rightarrow}\Psi_{t}(g)$. Suppose first that $g$ is bounded. Then
$\mathbb{E}(\rvert\Psi_{t}^{n}(g)-\Psi_{t}(g)\rvert)\leq\Delta_{n}+\int_{0}^{t}\mathbb{E}(\rvert g(\hat{a}(s))-g(a(s))\rvert)\mathrm{d}s.$
Theorem 1 in [17] along with the boundedness of $g$ allow us to apply the Dominated Convergence twice to conclude that the right-hand side of the previous inequality goes to 0 as $n\rightarrow\infty$. Now suppose that $g$ is of polynomial growth of order $q>0$ (when $q=0$ $g$ is bounded). Let $\psi:\mathbb{R}^{+}\rightarrow[0,1]$ be a function of class $\mathcal{C}^{\infty}$ such that $\mathbf{1}_{[1,\infty)}\leq\psi\leq\mathbf{1}_{[1/2,\infty)}$, and for every $m\in\mathbb{N}$, set $\psi_{m}(x):=1-\psi(\rvert x\rvert/m)$ as well as $\tilde{\psi}_{m}(x)=\psi(\rvert x\rvert/m).$ With this notation, we decompose
$\Psi_{t}^{n}(g)=\Psi_{t}^{n}(g_{m})+\Psi_{t}^{n}(\tilde{g}_{m}),$
where $g_{m}=g\psi_{m}$ and $\tilde{g}_{m}=g\tilde{\psi}_{m}$. Since for each $m$, $g_{m}$ is continuous and bounded, the first part of the proof gives that $\Psi_{t}^{n}(g_{m})\overset{\mathbb{P}}{\rightarrow}\int_{0}^{t}g_{m}(a(s))\mathrm{d}s$. In view that $\int_{0}^{t}g_{m}(a(s))\mathrm{d}s\rightarrow\Psi_{t}(g),$ as $m\rightarrow\infty$, we are left to show that for every $\eta>0$
$\lim_{m\rightarrow\infty}\limsup_{n\rightarrow\infty}\mathbb{P}(\rvert\Psi_{t}^{n}(\tilde{g}_{m})\rvert>\eta)=0.$ (43)
Since $g$ is of polynomial growth of order $q>0$, we have that for every $m\geq 1$, $\rvert\tilde{g}_{m}(x)\rvert\leq C_{q}\rvert x\rvert^{q}\mathbf{1}_{\rvert x\rvert>m}$, by the definition of $\psi$. Therefore, (43) will be achieved if
$\lim_{m\rightarrow\infty}\limsup_{n\rightarrow\infty}\mathbb{P}(\rvert Y_{n}^{m}(t)\rvert>\eta)=0,$ (44)
where
$Y_{n}^{m}(t)=\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}\rvert\hat{a}(l\Delta_{n})\rvert^{q}\mathbf{1}_{\{\rvert\hat{a}(l\Delta_{n})\rvert>m\}}.$
Relation (25) implies that
$Y_{n}^{m}(t)$ $\lesssim_{q}\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}\rvert\tilde{a}(l\Delta_{n})\rvert^{q}\mathbf{1}_{\{\rvert\hat{a}(l\Delta_{n})\rvert>m\}}$
$+\rvert\frac{1}{n}S_{n}\rvert^{q}\frac{1}{(n\Delta_{n})^{p}}\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}\rvert X_{n\Delta_{n}}-X_{l\Delta_{n}}\rvert^{q}$
$=:C_{q}(Y_{n}^{m,1}+Y_{n}^{2}).$
As in the proof of Lemma 4, we see that $Y_{n}^{2}\overset{\mathbb{P}}{\rightarrow}0$. Hence,
$\limsup_{n\rightarrow\infty}\mathbb{P}(\rvert Y_{n}^{m}(t)\rvert>\eta)\leq\limsup_{n\rightarrow\infty}\mathbb{P}(C_{q}\rvert Y_{n}^{m,1}\rvert>\eta/2).$ (45)
Moreover, Lemma 3 and Jensen’s inequality guarantee that
$\mathbb{E}(\rvert Y_{n}^{m,1}\rvert)\lesssim_{q}$ $\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}\mathbb{E}(\rvert\tilde{a}(l\Delta_{n})-\mathbb{E}[\tilde{a}(l\Delta_{n})]\rvert^{q}\mathbf{1}_{\{\rvert\hat{a}(l\Delta_{n})\rvert>m\}})$
$+\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}\mathbb{E}[\tilde{a}(l\Delta_{n})]^{q}\mathbb{P}(\rvert\hat{a}(l\Delta_{n})\rvert>m)$
$\lesssim$ $\frac{t}{(n\Delta_{n})^{q/2}}+\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}a(s)^{q}\mathrm{d}s\mathbb{P}(\rvert\hat{a}(l\Delta_{n})\rvert>m).$
By Markov’s and Hölder’s inequalities, we further have that
$\mathbb{P}(\rvert\hat{a}(l\Delta_{n})\rvert>m)\leq$ $\mathbb{P}(\rvert\tilde{a}(l\Delta_{n})\rvert>m/2)+\mathbb{P}(\rvert\frac{1}{n\Delta_{n}}(X_{n\Delta_{n}}-X_{l\Delta_{n}})\frac{1}{n}S_{n}\rvert>m/2)$
$\lesssim$ $\frac{1}{m}\mathbb{E}(\rvert\tilde{a}(l\Delta_{n})\rvert)+\frac{1}{m}\frac{1}{n\Delta_{n}}\mathbb{E}(\rvert X_{n\Delta_{n}}-X_{l\Delta_{n}}\rvert^{2})^{1/2}\mathbb{E}(\rvert\frac{1}{n}S_{n}\rvert^{2})^{1/2}.$
Therefore,
$\mathbb{E}(\rvert Y_{n}^{m,1}\rvert)\lesssim\frac{1}{(n\Delta_{n})^{q/2}}+\frac{1}{m}\left(1+\frac{1}{n\Delta_{n}}\mathbb{E}(\rvert\frac{1}{n}S_{n}\rvert^{2})^{1/2}\right).$
Relation (44) now follows from (45), Markov’s inequality, and the previous estimate.∎
Our next goal is to unify the proof of Theorems 2 and 4. To do so, we use $\mathcal{S}^{n}(g)$ ($\mathcal{S}(g)$) to denote either $\Psi^{n}(g)$, $\Lambda^{n}(g)$, or $\bar{\Lambda}^{n}(g)$ ($\Psi(g)$, $\Lambda(g)$, or $\bar{\Lambda}(g)$) and note that
$\mathcal{S}_{t}^{n}(g)=\Delta_{n}\sum_{l=0}^{n-1}c_{l}^{n}(t)g(\hat{a}(l\Delta_{n})),\,\,\,0\leq t\leq T,$ (46)
where
$c_{l}^{n}(t)=\begin{cases}\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}&\text{if }\mathcal{S}^{n}(g)=\Psi^{n}(g);\\ \mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}&\text{if }\mathcal{S}^{n}(g)=\Lambda^{n}(g);\\ \mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}&\text{if }\mathcal{S}^{n}(g)=\bar{\Lambda}^{n}(g).\end{cases}$ (47)
Then, if $g$ is of class $\mathcal{C}^{1}$, we deduce from the Mean-Value Theorem that there is $\hat{a}(l\Delta_{n})\land a(l\Delta_{n})\leq\theta_{l}\leq\hat{a}(l\Delta_{n})\lor a(l\Delta_{n}),$ such that
$\mathcal{S}_{t}^{n}(g)-\mathcal{S}_{t}(g)=U_{t}^{n}+\mathcal{R}_{t}^{n},\,\,\,\forall\,0\leq t\leq T,$
where
$U_{t}^{n}=\Delta_{n}\sum_{l=0}^{n-1}c_{l}^{n}(t)\dot{g}(\theta_{l})\left[\hat{a}(l\Delta_{n})-a(l\Delta_{n})\right],$ (48)
and
$\mathcal{R}_{t}^{n}=\Delta_{n}\sum_{l=0}^{n-1}c_{l}^{n}(t)g(a(l\Delta_{n}))-\mathcal{S}_{t}(g).$ (49)
Therefore, the conclusion of Theorems 2 and 4 will follow if the processes $U^{n}$ and $\mathcal{R}^{n}$ are asymptotically negligible. Please note that below we will always write $m_{n}=n$ or $m_{n}=N_{n}$, depending whether $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$, respectively.
Proof of Theorems 2 and 4.
Let $g\in\mathfrak{C}_{p,q}^{1}$ with $p,q\geq 1$ and note that in this situation
$\ \mathcal{R}_{t}^{n}=$ $\int_{t}^{m_{n}\Delta_{n}}[g(a(\tau_{n}(s)))-g(a(s))]\mathrm{d}s+\int_{[t/\Delta_{n}]\Delta_{n}}^{t}g(a(\tau_{n}(s)))\mathrm{d}s$ (50)
$-\int_{m_{n}\Delta_{n}}^{\infty}g(a(s))\mathrm{d}s.$
From (11), the mapping $s\mapsto g(a(s))$ is continuous with $\rvert g(a(s))\rvert\leq C_{a(0)}a(s)$. Therefore, uniformly on $t\in[0,T]$
$\int_{[t/\Delta_{n}]\Delta_{n}}^{t}g(a(\tau_{n}(s)))\mathrm{d}s-\int_{m_{n}\Delta_{n}}^{\infty}g(a(s))ds\rightarrow 0.$
Similarly, we deduce
$\int_{t}^{m_{n}\Delta_{n}}\rvert g(a(\tau_{n}(s)))-g(a(s))\rvert\mathrm{d}s$ $\lesssim_{a}\int_{0}^{m_{n}\Delta_{n}}[a(\tau_{n}(s))-a(s)]\mathrm{d}s$
$\leq\int_{0}^{\Delta_{n}}[a(0)-a(s)]\mathrm{d}s+\int_{0}^{\Delta_{n}}a(s)\mathrm{d}s$
$-\int_{(N_{n}-1)\Delta_{n}}^{N_{n}\Delta_{n}}a(s)\mathrm{d}s,$
$\lesssim_{a}\Delta_{n}$
This shows that $\mathcal{R}^{n}\overset{u.c.p}{\rightarrow}0$. Finally, since $\rvert\dot{g}(x)\rvert\lesssim\rvert x\rvert^{p}\mathbf{1}_{\rvert x\rvert\leq 1}+\rvert x\rvert^{q}\mathbf{1}_{\rvert x\rvert>1}$, we conclude that
$\sup_{t\in[0,T]}\rvert U_{t}^{n}\rvert\lesssim$ $\Delta_{n}\sum_{l=0}^{m_{n}-1}\left(\rvert\hat{a}(l\Delta_{n})-a(l\Delta_{n})\rvert^{p\land q+1}+\rvert\hat{a}(l\Delta_{n})-a(l\Delta_{n})\rvert^{p\lor q+1}\right)$ (51)
$+\Delta_{n}\sum_{l=0}^{m_{n}-1}a(l\Delta_{n})^{p\land q}\rvert\hat{a}(l\Delta_{n})-a(l\Delta_{n})\rvert$
$=$ $\mathrm{O}_{\mathbb{P}}(\frac{m_{n}}{n}(n\Delta_{n})^{\frac{1-p\land q}{2}})+\mathrm{O}_{\mathbb{P}}(1/\sqrt{n\Delta_{n}})+\mathrm{o}(1).$
where in the last step we applied Lemma 4 and the fact that $p,q\geq 1$. If $p,q>1$ and $m_{n}=n$ (as in Theorem 2), the right-hand side of (51) is clearly asymptotically negligible. On the other hand if $p\land q=1$ and $m_{n}=N_{n}$ satisfies (12) we have that the right-hand side of (51) is of order $\mathrm{O}_{\mathbb{P}}(\frac{N_{n}}{n})$, which by assumption converges to 0 as $n\rightarrow\infty$.∎
We conclude this section by presenting a proof for Theorem 3.
Proof of Theorem 3.
Arguing as in the preceding proof, we have that
$\Delta_{n}\sum_{l=0}^{n-1}\hat{a}(l\Delta_{n})^{2}-\int_{0}^{\infty}a(s)^{2}\mathrm{d}s=B_{n}+\mathrm{o}_{\mathbb{P}}(1)$
where $B_{n}=\Delta_{n}\sum_{l=0}^{n-1}\left[\hat{a}(l\Delta_{n})-a(l\Delta_{n})\right]^{2}$. Let
$b_{l}^{n}:=\hat{a}(l\Delta_{n})-a(l\Delta_{n})-\left(\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))\right).$
Arguing exactly as in the proof of Lemma 4, we have that
$\mathcal{E}_{n}:=\Delta_{n}\sum_{l=0}^{n-1}(b_{l}^{n})^{2}\overset{\mathbb{P}}{\rightarrow}0.$
Thus, by setting $B_{n}^{\prime}:=\Delta_{n}\sum_{l=0}^{n-1}\left[\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))\right]^{2}$ and applying the reverse triangle inequality we deduce that
$\left|\sqrt{B_{n}}-\sqrt{B_{n}^{\prime}}\right|\leq\sqrt{\mathcal{E}_{n}}\overset{\mathbb{P}}{\rightarrow}0.$
Therefore we only need to show that $B_{n}^{\prime}\overset{\mathbb{P}}{\rightarrow}B$. Now, using the random variables defined in (33), we set
$B_{n}^{\prime\prime}:=\Delta_{n}\sum_{l=0}^{n-1}\left[\frac{1}{n\Delta_{n}}\sum_{j=l}^{n}(\xi_{j,l,n}^{\prime(1)}+\xi_{j,l,n}^{(3)})\right]^{2},$
where $\xi_{j,l,n}^{\prime(1)}=\chi_{j}\beta_{j,j-1,l}^{(2)}$ as well as put (recall the decomposition in (32))
$\mathcal{E}_{n}^{\prime}:=\Delta_{n}\sum_{l=1}^{n-1}\left[\frac{1}{n\Delta_{n}}\sum_{j=l}^{n-1}\left(R_{j,l,n}+\xi_{j,l,n}^{(1)}+\xi_{j,l,n}^{(2)}+\xi_{j,l,n}^{(4)}\right)\right]^{2}.$
Now, a simple application of equation (63) in Appendix A and the definition of $R_{j,l,n}$, let us conclude that $\mathbb{E}(\mathcal{E}_{n}^{\prime})\rightarrow 0.$ This shows, just as above, that the limit of $B_{n}^{\prime\prime}$ is the same as the one of $B_{n}^{\prime}$ (ergo of $B_{n}$). Similar arguments used in the proof of Lemma 7 in A allow us to conclude that
$B_{n}^{\prime\prime}$ $:=\frac{\Delta_{n}}{(n\Delta_{n})^{2}}\sum_{j=1}^{n-1}\sum_{l=1}^{j}\left[(\xi_{j,l,n}^{\prime(1)})^{2}+(\xi_{j,l,n}^{(3)})^{2}+2\xi_{j,l,n}^{\prime(1)}\xi_{j,l,n}^{(3)}\right]$
$+\frac{2\Delta_{n}}{(n\Delta_{n})^{2}}\sum_{j=2}^{n-1}\sum_{l=1}^{j-1}(\xi_{j,l,n}^{\prime(1)}+\xi_{j,l,n}^{(3)})\sum_{j^{\prime}=l}^{j-1}(\xi_{j^{\prime},l,n}^{\prime(1)}+\xi_{j^{\prime},l,n}^{(3)})$
$=\frac{\Delta_{n}}{(n\Delta_{n})^{2}}\sum_{j=1}^{n-1}\sum_{l=1}^{j}\mathbb{E}\left[(\xi_{j,l,n}^{\prime(1)})^{2}+(\xi_{j,l,n}^{(3)})^{2}+2\xi_{j,l,n}^{\prime(1)}\xi_{j,l,n}^{(3)}\right]+\mathrm{o}_{\mathbb{P}}(1).$
The stated convergence now follows by invoking once again (63) in A.∎
4.3 Proof of Theorems 5 and 6
Similarly as in the previous subsection we fix $T>0$ and choose $n$ large enough such that $(m_{n}-1)\Delta_{n}>T$, where, as before $m_{n}=n$ or $m_{n}=N_{n}$, depending whether $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$, respectively. Once again we will unify our proofs by means of (46). Specifically, for every function $g$ of class $\mathcal{C}^{2}$ we have that
$Z_{t}^{n}:=\sqrt{n\Delta_{n}}\left(\mathcal{S}_{t}^{n}(g)-\mathcal{S}_{t}^{n}(g)\right)=V_{t}^{n}+\sum_{\ell=1}^{5}U_{t}^{n,\ell},$
where (see (47) and (49))
$V_{t}^{n}:=\sqrt{n\Delta_{n}}\Delta_{n}\sum_{l=0}^{n-1}c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))\left[\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))\right],$ (52)
and
$U_{t}^{n,1}=$ $\frac{\sqrt{n\Delta_{n}}}{2}\Delta_{n}\sum_{l=0}^{n-1}c_{l}^{n}(t)g^{\prime\prime}(\theta_{l})\left[\hat{a}(l\Delta_{n})-a(l\Delta_{n})\right]^{2};$
$U_{t}^{n,2}=$ $\sqrt{n\Delta_{n}}\sum_{l=0}^{n-1}c_{l}^{n}(t)\frac{(n-l)}{n}\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}\dot{g}(a(l\Delta_{n}))[a(s)-a(l\Delta_{n})]\mathrm{d}s;$
$U_{t}^{n,3}=$ $-\sqrt{\frac{\Delta_{n}}{n}}\sum_{l=0}^{n-1}c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))a(l\Delta_{n})l\Delta_{n};$
$U_{t}^{n,4}=$ $\frac{1}{n}S_{n}\sqrt{\frac{\Delta_{n}}{n}}\sum_{l=0}^{n-1}c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))(X_{n\Delta_{n}}-X_{l\Delta_{n}});$
$U_{t}^{n,5}=$ $\sqrt{n\Delta_{n}}\mathcal{R}_{t}^{n}.$
The next result shows that within our framework, the leading term is $V^{n}$.
Lemma 5.
Under the assumptions of Theorems 5 and 6, $U^{n,i}\overset{u.c.p}{\rightarrow}0$ for $i=1,\ldots,5$.
Proof.
That $U_{t}^{n,1}$ is negligible when $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}$(i.e. when $\mathcal{S}^{n}(g)=\Psi^{n}(g)$) can be shown exactly in the same way as in the proof of Theorems 2 and 4 (just replace $m_{n}$ by $T/\Delta_{n}]$ in (51)). Now, if $g\in\mathfrak{C}_{p,q}^{2}$ and either $m_{n}=n$ or $m_{n}=N_{n}$, we deduce as in (51) that uniformly on $t\in[0,T]$
$U_{t}^{n,1}=\mathrm{O}_{\mathbb{P}}(\frac{m_{n}}{n}(n\Delta_{n})^{\frac{1-p\land q}{2}})+\mathrm{O}_{\mathbb{P}}(\int_{0}^{m_{n}\Delta_{n}}a(s)^{p\land q}\mathrm{d}s/\sqrt{n\Delta_{n}})+\mathrm{o}(1).$ (53)
Since $m_{n}/n$ is bounded, we have that $U^{n,1}\overset{u.c.p}{\longrightarrow}0$ when either $p,q>1$ or $p\land q=1$ and $m_{n}=N_{n}$. If instead $0\leq p\land q<1$, suppose that $m_{n}=N_{n}$ and that Assumption 2 holds with $\kappa$ satisfying (16). To check that in this situation the right-hand side of (53) is still negligible as $n\rightarrow\infty$, it is enough to see that $\sqrt{n\Delta_{n}}\frac{N_{n}}{n}\rightarrow 0.$ If $\kappa\leq 1/2$ the result is obvious, so suppose that $1/2<\kappa<\frac{1}{2}(1+\frac{1}{\varpi}).$ Then by noticing that $\frac{1}{1-2(1-\kappa)}>\varpi$, we obtain that
$\sqrt{n\Delta_{n}}\frac{N_{n}}{n}\leq C(n\Delta_{n}^{\frac{1}{1-2(1-\kappa)}})^{\frac{1-2(1-\kappa)}{2}}\rightarrow 0.$
Next, let $\tilde{m}_{n}=[T/\Delta_{n}]$ if $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}$ and $\tilde{m}_{n}=m_{n}$ otherwise and note that Assumption 1 implies that the estimate in (42) is valid. Thus,
$\sup_{t\in[0,T]}\rvert U_{t}^{n,2}\rvert\lesssim\sqrt{n\Delta_{n}^{3}}\int_{0}^{\tilde{m}_{n}\Delta_{n}}\dot{g}(a(\tau_{n}(s)))\phi(s)\mathrm{d}s.$
Since $\sqrt{n\Delta_{n}^{3}}\rightarrow 0,$ and $\phi$ is bounded ($\phi$ is càdlàg and vanishes at $\infty$), to see that the result holds for $i=2,$ we only need to verify that under our assumptions the integral $\int_{0}^{m_{n}\Delta_{n}}\dot{g}[a(\tau_{n}(s))]\mathrm{d}s$ is bounded over $n$. Indeed, if $\tilde{m}_{n}=[T/\Delta_{n}]$ the statement follows trivially by the continuity of $\dot{g}$. If $\tilde{m}_{n}=m_{n}$ and $g\in\mathfrak{C}_{p,q}^{2}$, Remark 1 implies that
$\int_{0}^{m_{n}\Delta_{n}}\rvert\dot{g}[a(\tau_{n}(s))]\rvert\mathrm{d}s\lesssim\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))^{1+p}\mathrm{d}s\lesssim\Delta_{n}+\int_{0}^{\infty}a(s)^{1+p}\mathrm{d}s,$ (54)
just as required. For $i=3$, we observe that (recall that $xa(x)\rightarrow 0$)
$\sup_{t\in[0,T]}\rvert U_{t}^{n,3}\rvert\lesssim_{a}\frac{1}{\sqrt{n\Delta_{n}}}\int_{0}^{\tilde{m}_{n}\Delta_{n}}\dot{g}[a(\tau_{n}(s))]\mathrm{d}s\rightarrow 0,$
due to (54). Similar arguments let us infer that $\sup_{t\in[0,T]}\rvert U_{t}^{n,4}\rvert\overset{\mathbb{P}}{\rightarrow}0$. To deal with the final case suppose first that $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}$. Thus, in view that $g$ of class $\mathcal{C}^{2}$ and $a$ is Lipschitz (thanks to Assumption 1), we deduce that for all $t\in[0,T]$
$\sqrt{n\Delta_{n}}\rvert\mathcal{R}_{t}^{n}\rvert$ $\lesssim\sqrt{n\Delta_{n}}\int_{0}^{[t/\Delta_{n}]\Delta_{n}}\left|g(a(\tau_{n}(s)))-g(a(s))\right|\mathrm{d}s+\sqrt{n\Delta_{n}^{3}}$
$\lesssim\sqrt{n\Delta_{n}^{3}}\rightarrow 0.$
Similarly, if either $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$ and $g\in\mathfrak{C}_{p,q}^{2},$ we get from Remark 1 that
$\rvert\mathcal{R}_{t}^{n}\rvert\lesssim\int_{0}^{m_{n}\Delta_{n}}\rvert g(a(\tau_{n}(s)))-g(a(s))\rvert\mathrm{d}s+\int_{m_{n}\Delta_{n}}^{\infty}\rvert g(a(s))\rvert\mathrm{d}s+\Delta_{n}.$
Moreover, from (11) we deduce that $\rvert g(a(s))\rvert\leq Ca(s)^{2+p}$ for all $s>0$ and by the Mean-Value Theorem and the Lipshchitz condition, there is $x_{s,n}\in[a(s),a(\tau_{n}(s))]$, such that
$\rvert g(a(\tau_{n}(s)))-g(a(s))\rvert\lesssim\rvert\dot{g}(x_{s,n})\rvert\Delta_{n}\lesssim\Delta_{n}a(\tau_{n}(s))^{p+1},$
where in the second inequality we used also (11). A further application of Assumption 1 and (54) along with the previous estimates allow us to conclude that for $n$ large, uniformly on $t\in[0,T]$
$\sqrt{n\Delta_{n}}\rvert\mathcal{R}_{t}^{n}\rvert=\mathrm{o}(1)+\sqrt{n\Delta_{n}}(m_{n}\Delta_{n})^{1-\alpha(p+2)}.$
Therefore, if $m_{n}=n$, we have $\sqrt{n\Delta_{n}}(m_{n}\Delta_{n})^{1-\alpha(p+2)}=(n\Delta_{n})^{\frac{3}{2}-\alpha(p+2)}\rightarrow 0$ because $\alpha>1$. In contrast, if $m_{n}=N_{n}$ and Assumption 2 holds with $\kappa$ as in (16), we have that
$\sqrt{n\Delta_{n}}(m_{n}\Delta_{n})^{1-\alpha(p+2)}\leq C(n\Delta_{n}^{\frac{2((2+p)\alpha-1)-1}{2\kappa((2+p)\alpha-1)-1}})^{-\frac{2\kappa((2+p)\alpha-1)-1}{2}}\rightarrow 0,$
due to the fact that $\kappa>\frac{1}{\varpi}+\frac{\varpi-1}{2\varpi((2+p)\alpha-1)}>\frac{1}{2((2+p)\alpha-1)}.$ This concludes the proof.∎
The next step is to write $V^{n}$ defined in (52) as a sum of of martingale differences plus an error term. Once we are done with this, the proof of Theorems 5 and 6 will follow as an application of e.g. Theorem 6.1 in [8] (c.f. Theorem IX 7.19 in [12]). We do this by means of the decomposition (32):
$V_{t}^{n}=V_{t}^{\prime n}+U_{t}^{\prime,n},$
where
$V_{t}^{\prime n}$= $\sqrt{\frac{\Delta_{n}}{n}}\sum_{l=0}^{n-1}\sum_{j=l}^{n-1}c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))\varXi_{j,l,n};$ (55)
$U_{t}^{\prime,n}$= $\sqrt{\frac{\Delta_{n}}{n}}\sum_{l=0}^{n-1}\sum_{j=l}^{n-1}c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))R_{j,l,n}.$
Clearly $\varXi_{j,l,n}$ is $\mathscr{G}_{j}^{n}$-measurable (see (23)) and by construction $\mathbb{E}(\varXi_{j,l,n}\rvert\mathscr{G}_{j-1}^{n})=0$, so is a $(\mathscr{G}_{j}^{n})_{j\geq 0}$ martingale difference. Let us now see that $U^{\prime,n}$ is an error term.
Lemma 6.
Under the assumptions of Theorems 5 and 6, $U_{t}^{\prime,n}\overset{\mathbb{P}}{\rightarrow}0$.
Proof.
As before we may and do assume that $L$ has mean 0. Since each $\rho_{j,l,n}^{(\ell)}$ (see (34)) is a martingale difference, we only need to show that for $\ell=1,2,3$
$\frac{1}{n\Delta_{n}}\mathcal{E}_{n}^{(\ell)}(t):=\frac{1}{n\Delta_{n}}\sum_{j=0}^{n}\mathbb{E}\left[\left(\Delta_{n}\sum_{l=0}^{j}c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))\rho_{j,l,n}^{(\ell)}\right)^{2}\right]\rightarrow 0.$
To see that this is certainly the case, first note that $\Delta_{n}\sum_{l=0}^{j}\rvert c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))\rvert$ is uniformly bounded. Indeed, if $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}$, by the continuity of $\dot{g}$ and $a$
$\Delta_{n}\sum_{l=0}^{j}\rvert c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))\rvert\leq\Delta_{n}\sum_{l=0}^{[T/\Delta_{n}]}\lvert\dot{g}(a(l\Delta_{n}))\rvert\leq CT.$
If instead either $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$ and $g\in\mathfrak{C}_{p,q}^{2},$ we deduce from (54) that our statement is also true in this case. Hence, by Jensen’s inequality
$\mathcal{E}_{n}^{(\ell)}(t)\lesssim\Delta_{n}\sum_{j=0}^{n}\sum_{l=0}^{j}\rvert c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))\rvert\mathbb{E}[(\rho_{j,l,n}^{(\ell)})^{2}].$
Furthermore, by Lemma 1
$\mathbb{E}[(\rho_{j,l,n}^{(\ell)})^{2}]$ $\leq a(0)\Delta_{n}\int_{(n-j)\Delta_{n}}^{\infty}a(s)\mathrm{d}s;\,\ell=1,3;$
$\mathbb{E}[(\rho_{j,l,n}^{(2)})^{2}]$ $\leq\mathbb{E}[X_{0}^{2}]\int_{j\Delta_{n}}^{(j+1)\Delta_{n}}a(n\Delta_{n}-s)\mathrm{d}s.$
Therefore
$\frac{1}{n\Delta_{n}}\mathcal{E}_{n}^{(2)}(t)\lesssim\frac{1}{n\Delta_{n}}\sum_{j=0}^{n-1}\int_{j\Delta_{n}}^{(j+1)\Delta_{n}}a(n\Delta_{n}-s)\mathrm{d}s\leq\mathrm{Leb}(A)/n\Delta_{n}\rightarrow 0,$
and by the Dominated Convergence Theorem we also have that
$\frac{1}{n\Delta_{n}}\mathcal{E}_{n}^{(\ell)}(t)\leq C\frac{1}{\Delta_{n}n}\int_{0}^{n\Delta_{n}}\int_{r}^{\infty}a(s)\mathrm{d}s\mathrm{d}r\rightarrow 0,\,\,\ell=1,3,$
as required. ∎
We are now ready to present a proof for the CLT associated to our statistics. Note that in view of the previous two lemmas and the definition of $\tilde{a}$ (see (26)), we may and do assume that $L$ has mean 0 for the rest of this work.
Proof of Theorems 5 and 6.
We start making some observations. As pointed out above, we only need to concentrate on $V^{n}$ (see (52)). Note also that thanks to Lemma 3, for all $0\leq s\leq t\leq T$ we have
$\mathbb{E}(\rvert V_{t}^{n}-V_{s}^{n}\rvert^{2})^{1/2}\leq C\Delta_{n}\sum_{l=[s/\Delta_{n}]}^{[t/\Delta_{n}]-1}\rvert g(a(l\Delta_{n}))\rvert\leq C\Delta_{n}([t/\Delta_{n}]-[s/\Delta_{n}]),$
Therefore, according to [5] (Theorem 13.5 and equation (13.14)), Remark 2, and Lemma 6, we only need to show that the finite-dimensional distributions of $V^{\prime n}$ (for its definition see (55)) converge to those of $Z$, with $Z$ as in the theorems. We will use the Cramér-Wold methodology. Fix $M\in\mathbb{N}$ and let $0=t_{0}\leq t_{1}\leq\cdots\leq t_{M}\leq T$ and $z_{1},z_{2},\ldots,z_{M}\in\mathbb{R}$. By setting $\overline{c}_{l}^{n}=\sum_{k=1}^{M}z_{k}c_{l}^{n}(t_{k})\dot{g}(a(l\Delta_{n})),$ and
$\zeta_{j}^{n}=\Delta_{n}\sum_{l=0}^{j}\overline{c}_{l}^{n}\varXi_{j,l,n},$ (56)
in which $\varXi_{j,l,n}=\sum_{\ell=1}^{4}\xi_{j,l,n}^{(\ell)}$ (see (33)) we have that
$\sum_{k=1}^{M}z_{k}V_{t_{k}}^{\prime n}=\frac{1}{\sqrt{n\Delta_{n}}}\sum_{j=0}^{n}\zeta_{j}^{n}.$
As pointed out above, $\zeta_{j}^{n}$ is a $(\mathscr{G}_{j}^{n})_{j\geq 0}$-martingale difference. Therefore, according to Theorem 6.1 in [8] (c.f. Theorem IX 7.19 in [12]), we need to show that for some $r>2$
$1.\,\sigma_{n}^{2}:=\frac{1}{n\Delta_{n}}\sum_{j=1}^{n}\mathbb{E}[(\zeta_{j}^{n})^{2}\mid\mathscr{G}_{j-1}^{n}]$ $\overset{\mathbb{P}}{\rightarrow}\sum_{k,k^{\prime}}z_{k}z_{k^{\prime}}\mathbb{E}(Z_{t_{k}}Z_{t_{k^{\prime}}}),$ (57)
$2.\,\frac{1}{(n\Delta_{n})^{r/2}}\sum_{j=1}^{n}\mathbb{E}(\mid\zeta_{j}^{n}\mid^{r})$ $\overset{\mathbb{P}}{\rightarrow}0.$ (58)
We verify relations (57) and (58) as follow:
Asymptotic Variance: In Lemmas 7 and 8 in Section A, it is shown that
$\sigma_{n}^{2}=\Delta_{n}^{2}\sum_{l=0}^{n-1}\sum_{l^{\prime}=0}^{n-1}\overline{c}_{l}^{n}\overline{c}_{l^{\prime}}^{n}\left(\sum_{\ell,\ell^{\prime}=1}^{4}F_{n}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})\right)+\mathrm{o}_{\mathbb{P}}(1),$ (59)
where $F_{n}^{(\ell,\ell^{\prime})}$ is a sequence of bounded functions vanishing outside of $0,n\Delta_{n}]$ and such that (recall that $\tau_{n}(s)=[s/\Delta_{n}]\Delta_{n}$) $\sum_{\ell,\ell^{\prime}=1}^{4}F_{n}^{(\ell,\ell^{\prime})}(\tau_{n}(s),\tau_{n}(r))\rightarrow\Sigma_{a}(s,r)$, as $n\rightarrow\infty$. This means that if $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$, then as $n\rightarrow\infty$, the continuity of $\dot{g}$ implies that
$\sigma_{n}^{2}\overset{\mathbb{P}}{\rightarrow}\sum_{\ell,\ell^{\prime}=1}^{4}\sum_{q,q^{\prime}=1}^{M}z_{q}z_{q^{\prime}}\int_{0}^{t_{q}}\int_{0}^{t_{q^{\prime}}}\dot{g}(a(s))\dot{g}(a(r))\Sigma_{a}(s,r)\mathrm{d}r\mathrm{d}s,$
which is exactly (57). On the other hand, if either $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$
$\sigma_{n}^{2}=\sum_{\ell,\ell^{\prime}=1}^{4}\sum_{q,q^{\prime}=1}^{M}z_{q}z_{q^{\prime}}\int_{t_{q}}^{m_{n}\Delta_{n}}\int_{t_{q^{\prime}}}^{m_{n}\Delta_{n}}\dot{g}(a(\tau_{n}(s)))\dot{g}(a(\tau_{n}(r)))F_{n}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})\mathrm{d}r\mathrm{d}s,$ (60)
plus an error term that is asymptotically negligible. Therefore, in order to achieve (57) in this situation, it is enough to check that the (Generalized) Dominated Convergence Theorem (DCT from now on) can be applied in (60). To see that the latter is indeed possible, observe that from (11) we have that
$\rvert\dot{g}(a(\tau_{n}(s)))\rvert\lesssim\mathbf{1}_{s\leq\Delta_{n}}+a(s-\Delta_{n})\mathbf{1}_{s>\Delta_{n}}=:f_{n}(s)\rightarrow a(s),$ (61)
for all $s>0$. By the integrability and boundedness of $a$ we further have
$\int_{t}^{m_{n}\Delta_{n}}f_{n}(s)\mathrm{d}s=\int_{t}^{m_{n}\Delta_{n}}a(s)\mathrm{d}s+\mathrm{o}(1)\rightarrow\int_{t}^{\infty}a(s)\mathrm{d}s,\,\,\,t\geq 0.$ (62)
The previous two properties together with the boundedness of $F_{n}^{(\ell,\ell^{\prime})}$ justify the use of the Generalized DCT in (60).
Lyapunov’s condition: Arguing as in the proof of Lemma 6 we get that $\Delta_{n}\sum_{l=0}^{j}\rvert\overline{c}_{l}^{n}\rvert$ is uniformly bounded. In consequence, if we choose $2<r<\frac{r_{0}}{2}\land 2\alpha$, we infer from Jensen’s inequality, Lemma 2, and Assumption 1 that for every $\ell=1,2,3,4$
$\mathbb{E}(\mid\zeta_{j}^{n,(\ell)}\mid^{r})\leq\left(4\Delta_{n}\sum_{l=0}^{j}\rvert\overline{c}_{l}^{n}\rvert\right)^{r-1}\Delta_{n}\sum_{l=0}^{j}\rvert\overline{c}_{l}^{n}\rvert\mathbb{E}(\mid\xi_{j,l,n}^{(\ell)}\mid^{r})\leq C\Delta_{n},$
from which (58) trivially holds.∎
Appendix A Computing the AVAR
In this subsection we show that the approximation (59) is valid. We will constantly use the estimates and identities presented in Lemma 1 as well as the basic fact that the variance of a sum of martingale differences is the sum of each individual variance. Thus, the reader must be aware that we will often omit referring to these results unless the context is not clear enough. Recall that $m_{n}=n$ or $m_{n}=N_{n}$, depending whether $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$.
Lemma 7.
Let $\zeta_{j}^{n}=\sum_{\ell=1}^{4}\zeta_{j}^{n,(\ell)}$ with $\zeta_{j}^{n,(\ell)}$ as in (56). Under the assumptions of Theorems 5 and 6, it holds that
$\frac{1}{n\Delta_{n}}\sum_{j=0}^{n}\mathbb{E}[(\zeta_{j}^{n})^{2}\mid\mathscr{G}_{j-1}^{n}]=\frac{1}{n\Delta_{n}}\sum_{j=0}^{n}\mathbb{E}(\zeta_{j,n}^{2})+\mathrm{o}_{\mathbb{P}}(1).$
Proof.
Plainly
$\frac{1}{n\Delta_{n}}\sum_{j=0}^{n}\mathbb{E}[(\zeta_{j}^{n})^{2}\mid\mathscr{G}_{j-1}^{n}]=\sum_{\ell,\ell^{\prime}=1}^{4}\Delta_{n}^{2}\sum_{l=0}^{n-1}\sum_{l^{\prime}=0}^{n-1}\overline{c}_{l}^{n}\overline{c}_{l^{\prime}}^{n}\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}}^{n}\mathbb{E}(\xi_{j,l,n}^{(\ell)}\xi_{j,l^{\prime},n}^{(\ell^{\prime})}\mid\mathscr{G}_{j-1}^{n}).$
By the independent scattered property of $L$ we have that
$\mathbb{E}(\xi_{j,l,n}^{(1)}\xi_{j,l^{\prime},n}^{(1)}\mid\mathscr{G}_{j-1}^{n})=$ $\mathbb{E}(\chi_{j}^{2}\beta_{j,j,l}^{(1)}\beta_{j,j,l^{\prime}}^{(1)})+\beta_{j,j-1,l^{\prime}}^{(2)}\mathbb{E}(\chi_{j}^{2}\beta_{j,j,l}^{(1)})$ (63)
$+\beta_{j,j-1,l}^{(2)}\mathbb{E}(\chi_{j}^{2}\beta_{j,j,l^{\prime}}^{(1)})+\beta_{j,j-1,l}^{(2)}\beta_{j,j-1,l^{\prime}}^{(2)}\mathbb{E}(\chi_{j}^{2})$
$-\mathbb{E}(\chi_{j}\beta_{j,j,l}^{(2)})\mathbb{E}(\chi_{j}\beta_{j,j,l^{\prime}}^{(2)});$
$\mathbb{E}(\xi_{j,l,n}^{(2)}\xi_{j,l^{\prime},n}^{(2)}\mid\mathscr{G}_{j-1}^{n})=$ $\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}\chi_{k}\chi_{k^{\prime}}\mathbb{E}[\beta_{k,j,l}^{(1)}\beta_{k^{\prime},j,l^{\prime}}^{(1)}];$
$\mathbb{E}(\xi_{j,l,n}^{(3)}\xi_{j,l^{\prime},n}^{(3)}\mid\mathscr{G}_{j-1}^{n})=$ $\sum_{k=l\lor l^{\prime}}^{j-1}\beta_{k,j-1,l}^{(2)}\beta_{k,j-1,l^{\prime}}^{(2)}\mathbb{E}[\alpha_{k+1,j}^{2}];$
$\mathbb{E}(\xi_{j,l,n}^{(4)}\xi_{j,l^{\prime},n}^{(4)}\mid\mathscr{G}_{j-1}^{n})=$ $\sum_{k=l\lor l^{\prime}}^{j-1}\mathbb{E}(\alpha_{k+1,j}^{2})\mathbb{E}(\beta_{k,j,l}^{(1)}\beta_{k,j,l^{\prime}}^{(1)})$
$+\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}\varsigma_{k^{\prime},j}\varsigma_{k,j}\mathbb{E}(\beta_{k,j,l}^{(1)}\beta_{k^{\prime},j,l^{\prime}}^{(1)}),$
where $\varsigma_{k,j}:=\sum_{m=k+1}^{j-1}\alpha_{k+1,m}$. Furthermore,
$\mathbb{E}(\xi_{j,l,n}^{(1)}\xi_{j,l^{\prime},n}^{(2)}\mid\mathscr{G}_{j-1}^{n})=$ $\sum_{k=l^{\prime}}^{j-1}\chi_{k}\beta_{j,j-1,l}^{(2)}\mathbb{E}[(\beta_{k,j,l^{\prime}}^{(1)})^{2}]$ (64)
$+\sum_{k=l^{\prime}}^{j-1}\chi_{k}\mathbb{E}(\beta_{k,l^{\prime},j}^{(1)}\beta_{j,j,l}^{(1)}\chi_{j})$
$-\mathbb{E}(\xi_{j,l,n}^{(1)}\xi_{j,l^{\prime},n}^{(3)}\mid\mathscr{G}_{j-1}^{n})=$ $\sum_{k=l^{\prime}}^{j-1}\beta_{j,j-1,l}^{(2)}\beta_{k,j-1,l^{\prime}}^{(2)}\mathbb{E}(\alpha_{k+1,j}\chi_{j})$
$+\sum_{k=l^{\prime}}^{j-1}\beta_{k,j-1,l^{\prime}}^{(2)}\mathbb{E}(\alpha_{k+1,j}\chi_{j}\beta_{j,j,l}^{(1)});$
$-\mathbb{E}(\xi_{j,l,n}^{(1)}\xi_{j,l^{\prime},n}^{(4)}\mid\mathscr{G}_{j-1}^{n})=$ $\sum_{k=l^{\prime}}^{j-1}\mathbb{E}(\beta_{j,j,l}^{(1)}\chi_{j}\alpha_{k+1,j}\beta_{k,j,l^{\prime}}^{(1)})$
$+\sum_{k=l^{\prime}}^{j-1}\varsigma_{k,j}\mathbb{E}(\beta_{j,j,l}^{(1)}\chi_{j}\beta_{k,j,l^{\prime}}^{(1)})$
$+\sum_{k=l^{\prime}}^{j-1}\varsigma_{k,j}\beta_{j,j-1,l}^{(2)}\mathbb{E}(\chi_{j}\beta_{k,j,l^{\prime}}^{(1)});$
as well as
$-\mathbb{E}(\xi_{j,l,n}^{(2)}\xi_{j,l^{\prime},n}^{(3)}\mid\mathscr{G}_{j-1}^{n})=$ $\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}\chi_{k}\beta_{k^{\prime},j-1,l^{\prime}}^{(2)}\mathbb{E}[\alpha_{k^{\prime}+1,j}^{2}]\mathbf{1}_{k^{\prime}+1\leq k-l};$ (65)
$-\mathbb{E}(\xi_{j,l,n}^{(2)}\xi_{j,l^{\prime},n}^{(4)}\mid\mathscr{G}_{j-1}^{n})=$ $\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}\chi_{k}\varsigma_{k^{\prime},j}\mathbb{E}[\beta_{k,j,l}^{(1)}\beta_{k^{\prime},j,l^{\prime}}^{(1)}];$
$\mathbb{E}(\xi_{j,l,n}^{(3)}\xi_{j,l^{\prime},n}^{(4)}\mid\mathscr{G}_{j-1}^{n})=$ $\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}\beta_{k,j-1,l}^{(2)}\varsigma_{k^{\prime},j}\mathbb{E}[\alpha_{k+1,j}\beta_{k^{\prime},j,l^{\prime}}^{(1)}]\mathbf{1}_{k+1\leq k^{\prime}-l^{\prime}}.$
We analyze now separately the terms
$\mathcal{V}_{n}^{(\ell,\ell^{\prime})}:=\Delta_{n}^{2}\sum_{l=0}^{n-1}\sum_{l^{\prime}=0}^{n-1}\overline{c}_{l}^{n}\overline{c}_{l^{\prime}}^{n}\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}}^{n}\mathbb{E}(\xi_{j,l,n}^{(\ell)}\xi_{j,l^{\prime},n}^{(\ell^{\prime})}\mid\mathscr{G}_{j-1}^{n}).$ (66)
We often use the fact that that $\Delta_{n}\sum_{l=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert$ is uniformly bounded (this can be seen as in the proof of Lemma 6). Furthermore, we are going to verify that
$\mathcal{V}_{n}^{(\ell,\ell^{\prime})}=\mathbb{E}(\mathcal{V}_{n}^{(\ell,\ell^{\prime})})+\Delta_{n}^{2}\sum_{l=1}^{n-1}\sum_{l^{\prime}=0}^{n-1}\overline{c}_{l}^{n}\overline{c}_{l^{\prime}}^{n}\frac{\mathcal{S}_{l,l^{\prime}}^{n}}{n\Delta_{n}},$ (67)
where can be written as $\mathcal{S}_{l,l^{\prime}}^{n}=\sum_{\iota=1}^{\mathrm{N}}\mathcal{S}_{l,l^{\prime}}^{n,\iota}$, $\mathrm{N}\in\mathbb{N}$ fixed, and each summand satisfies either
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\iota}\rvert^{2})\lesssim n\Delta_{n}\upsilon_{n},$ (68)
uniformly on $l,l^{\prime}$ and with with $\upsilon_{n}=\mathrm{o}(n\Delta_{n})$; or
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\iota}\rvert^{2})\lesssim n\Delta_{n}[\Delta_{n}(l\lor l^{\prime})+C],$ (69)
for some constant $C\geq 0$. Note that these inequalities will imply the desired result. Indeed, by Jensen’s inequality we have that
$\mathbb{E}(\rvert\mathcal{V}_{n}^{(\ell,\ell^{\prime})}-\mathbb{E}(\mathcal{V}_{n}^{(\ell,\ell^{\prime})})\rvert^{2})\lesssim\sum_{\iota=1}^{\mathrm{N}}\Delta_{n}^{2}\sum_{l=0}^{n-1}\sum_{l^{\prime}=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert\rvert\overline{c}_{l^{\prime}}^{n}\rvert\frac{1}{(n\Delta_{n})^{2}}\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\iota}\rvert^{2}).$
Therefore, for all $\iota=1,\ldots,\mathrm{N}$, either
$\Delta_{n}^{2}\sum_{l=0}^{n-1}\sum_{l^{\prime}=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert\rvert\overline{c}_{l^{\prime}}^{n}\rvert\frac{1}{(n\Delta_{n})^{2}}\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\iota}\rvert^{2})\lesssim\upsilon_{n}/n\Delta_{n}\rightarrow 0,\,\,n\rightarrow\infty,$
or
$\Delta_{n}^{2}\sum_{l=0}^{n-1}\sum_{l^{\prime}=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert\rvert\overline{c}_{l^{\prime}}^{n}\rvert\frac{1}{(n\Delta_{n})^{2}}\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\iota}\rvert^{2})\lesssim\frac{1}{n}\sum_{l=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert t_{l}+1/n\Delta_{n}.$
In the latter case we further have that as $n\rightarrow\infty$
  1. 1.
    If $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1},$ by the continuity of $\dot{g}$
    $\frac{1}{n}\sum_{l^{\prime}=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert t_{l}\lesssim T/n\Delta_{n}\rightarrow 0.$
  2. 2.
    Otherwise, arguing as in (54), we deduce that
    $\frac{1}{n}\sum_{l^{\prime}=0}^{n-1}\rvert\overline{c}_{l}^{n}\rvert t_{l}\lesssim\mathrm{O}(1/n\Delta_{n})+\frac{m_{n}}{n}\times\frac{1}{m_{n}\Delta_{n}}\int_{0}^{m_{n}\Delta_{n}}a(s)^{p+1}s\mathrm{d}s\rightarrow 0,$
    thanks to Assumption 1 and (12).
Therefore, in any situation, it holds that $\mathbb{E}(\rvert\mathcal{V}_{n}^{(\ell,\ell^{\prime})}-\mathbb{E}(\mathcal{V}_{n}^{(\ell,\ell^{\prime})})\rvert^{2})\rightarrow 0$, as $n\rightarrow\infty$, which is exactly the conclusion of this Lemma.
For convenience, when $\ell=\ell^{\prime}$, we will refer to $\mathcal{V}_{n}^{(\ell,\ell^{\prime})}$ as the conditional variance, otherwise we will use the adjective conditional covariance. We would also like to emphasize that when we present the decomposition (67) we are always going to use the notation $\mathcal{S}_{l,l^{\prime}}^{n}$ and its definition will vary from case to case.
Conditional Variances
$\ell=1:$
Set $\mathfrak{K}_{3}=\int x^{3}\nu(dx)$ and write
$\mathcal{S}_{l,l^{\prime}}^{n}:=$ $\mathfrak{K}_{3}\int_{t_{l}}^{t_{l+1}}a(s)\mathrm{d}s\sum_{j=l\lor l^{\prime}}^{n-1}\beta_{j,j-1,l^{\prime}}^{(2)}$
$+\mathfrak{K}_{3}\int_{t_{l}}^{t_{l+1}}a(s)\mathrm{d}s\sum_{j=l\lor l^{\prime}}^{n-1}\beta_{j,j-1,l}^{(2)}$
$+\int_{0}^{\Delta_{n}}a(s)\mathrm{d}s\sum_{j=l\lor l^{\prime}}^{n-1}[\beta_{j,j-1,l}^{(2)}\beta_{j,j-1,l^{\prime}}^{(2)}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\beta_{j,j-1,l^{\prime}}^{(2)})]$
$=:$ $\sum_{\iota=1}^{3}\mathcal{S}_{l,l^{\prime}}^{n,\iota},$
so that (67) holds. Since the sequence $\beta_{j,j-1,l^{\prime}}^{(2)}$ is $l^{\prime}$-dependent and $\mathbb{E}[(\beta_{j,j-1,l^{\prime}}^{(2)})^{2}]\leq\mathrm{Leb}(A),$ we have that
$\mathbb{E}\left[\left(\sum_{j=l\lor l^{\prime}}^{n}\beta_{j,j-1,l^{\prime}}^{(2)}\right)^{2}\right]\leq Cnl^{\prime},$
which adds up to $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,1}\rvert^{2})\lesssim n\Delta_{n}(l^{\prime}\lor l\Delta_{n}+C).$ Similar arguments can be applied ($\beta_{j,j-1,l}^{(2)}\beta_{j,j-1,l^{\prime}}^{(2)}$ is $l\lor l^{\prime}$-dependent) to show that the same estimate holds for $\mathcal{S}_{l,l^{\prime}}^{n,2}$ and $\mathcal{S}_{l,l^{\prime}}^{n,3}$, which is enough.
$\ell=2:$
By interchanging the order of summation we may write
$\sum_{j=l\lor l^{\prime}}^{n}\mathbb{E}(\xi_{j,l,n}^{(2)}\xi_{j,l^{\prime},n}^{(2)}\mid\mathscr{G}_{j-1}^{n})=\sum_{k=l}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{n-1}\chi_{k}\chi_{k^{\prime}}\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k^{\prime}-l^{\prime})}\cup_{j=k\lor k^{\prime}+1}^{n-1}\mathcal{P}_{A}^{n}(i,j)).$
Thus, (67) holds with
$\mathcal{S}_{l,l^{\prime}}^{n}:=$ $\sum_{k=l}^{n-1}[\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})]b_{k,l,l^{\prime}}+\sum_{k=l}^{n-1}\chi_{k}\nu_{k,l^{\prime},l}$
$+\sum_{k^{\prime}=l^{\prime}}^{n-1}\chi_{k^{\prime}}\tilde{\nu}_{k,l,l^{\prime}}$
$=:$ $\sum_{\iota=1}^{3}\mathcal{S}_{l,l^{\prime}}^{n,\iota}$
where
$b_{k,l,l^{\prime}}:=$ $\sum_{k^{\prime}=l^{\prime}}^{n-1}\mathbf{1}_{k^{\prime}=k}\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k^{\prime}-l^{\prime})}\cup_{j=k+1}^{n-1}\mathcal{P}_{A}^{n}(i,j));$
$\nu_{k,l,l^{\prime}}:=$ $\sum_{k^{\prime}=l^{\prime}}^{n-1}\chi_{k^{\prime}}\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k^{\prime}-l^{\prime})}\cup_{j=k+1}^{n-1}\mathcal{P}_{A}^{n}(i,j))\mathbf{1}_{k^{\prime}\leq k-1};$
$\tilde{\nu}_{k,l,l^{\prime}}:=$ $\sum_{k=l}^{n-1}\chi_{k}\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k^{\prime}-l^{\prime})}\cup_{j=k^{\prime}+1}^{n-1}\mathcal{P}_{A}^{n}(i,j))\mathbf{1}_{k\leq k^{\prime}-1}$
Now note that $b_{k,l,l^{\prime}}\leq\mathrm{Leb}(A)$ and $\mathbb{E}(\rvert\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})\rvert^{2}]\lesssim\Delta_{n}$. Furthermore
$\mathbb{E}(\rvert\chi_{k}\nu_{k,l^{\prime},l}\rvert^{2})\lesssim$ $\Delta_{n}^{2}\sum_{k^{\prime}=l^{\prime}}^{n-1}\mathrm{Leb}(\cup_{i=0}^{k^{\prime}-l^{\prime}}\cup_{j\geq k}\mathcal{P}_{A}^{n}(i,j))^{2}\mathbf{1}_{k^{\prime}\leq k}$
$\lesssim$ $\Delta_{n}\int_{0}^{n\Delta_{n}}\left(\int_{s}^{\infty}a(u)\mathrm{d}u\right)^{2}\mathrm{d}s$
$\lesssim$ $\Delta_{n}(C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)}),$
where in the last inequality we also applied Assumption 1. Similarly, we deduce that that $\mathbb{E}(\rvert\chi_{k^{\prime}}\tilde{\nu}_{k,l,l^{\prime}}\rvert^{2})\lesssim\Delta_{n}(C_{\alpha}+(n\Delta_{n})^{1+2(1-\alpha)}).$ In view that $\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})]$, $\chi_{k}\nu_{k,l^{\prime},l}$ and $\chi_{k^{\prime}}\tilde{\nu}_{k,l,l^{\prime}}$ are all martingale differences, it follows that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,1}\rvert^{2})=\sum_{k=l}^{n-1}\mathbb{E}(\rvert\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})\rvert^{2})b_{k,l,l^{\prime}}^{2}\leq\mathrm{Leb}(A)^{2}n\Delta_{n}.$
Likewise, for $\iota=2,3$ we have that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\iota}\rvert^{2})\lesssim n\Delta_{n}(C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)}).$
Therefore, in this situation (68) is satisfied with $\upsilon_{n}=C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)}.$
$\ell=3:$
By symmetry we may write
$\mathcal{V}_{n}^{(3,3)}$ $=\mathbb{E}(\mathcal{V}_{n}^{(3,3)})+\Delta_{n}^{2}\sum_{l=0}^{n-1}(\overline{c}_{l}^{n})^{2}\mathcal{S}_{l,l}^{n}$
$+2\Delta_{n}^{2}\sum_{l=1}^{n-1}\sum_{l^{\prime}=0}^{l-1}\overline{c}_{l}^{n}\overline{c}_{l^{\prime}}^{n}\frac{\mathcal{S}_{l,l^{\prime}}^{n}}{n\Delta_{n}},$
where
$\mathcal{S}_{l,l^{\prime}}^{n}:=\sum_{j=l}^{n}\sum_{k=l}^{j-1}[\beta_{k,j-1,l}^{(2)}\beta_{k,j-1,l^{\prime}}^{(2)}-\mathbb{E}(\beta_{k,j-1,l}^{(2)}\beta_{k,j-1,l^{\prime}}^{(2)})]\mathbb{E}[\alpha_{k+1,j}^{2}].$
The goal of this part is to show that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n}\rvert^{2})\lesssim n\Delta_{n}(l\Delta_{n}+\mathrm{Leb}(A)).$ (70)
The $\sigma$-additivity of $L$ implies that
$\beta_{k,j-1,l}^{(2)}=\sum_{m=k-l}^{j-1}\beta_{k,m,l}^{(1)}.$ (71)
Applying this and performing several tedious changes of order of summation we conclude that $\mathcal{S}_{l,l^{\prime}}^{n}=\sum_{\imath=1}^{7}\mathcal{S}_{l,l^{\prime}}^{n,\imath}$, where
$\mathcal{S}_{l,l^{\prime}}^{n,1}:=$ $\sum_{m=0}^{n-1}\sum_{k=l\lor m+1}^{(n-1)\land(m+l)}[\beta_{k,m,l}^{(1)}\beta_{k,m,l^{\prime}}^{(1)}-\mathbb{E}(\beta_{k,m,l}^{(1)}\beta_{k,m,l^{\prime}}^{(1)})]\sum_{j=k+1}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$
$\mathcal{S}_{l,l^{\prime}}^{n,2}:=$ $\sum_{m=l-l^{\prime}}^{n-1}\sum_{m^{\prime}=(l-l^{\prime})\lor(m-l^{\prime})+1}^{m-1}\sum_{k=l\lor m+1}^{(n-1)\land(m^{\prime}+l^{\prime})}\beta_{k,m,l}^{(1)}\beta_{k,m^{\prime},l^{\prime}}^{(1)}\sum_{j=k+1}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$
$\mathcal{S}_{l,l^{\prime}}^{n,3}:=$ $\sum_{m=0}^{n-1}\sum_{k=l\lor m+1}^{(n-1)\land(m+l)}\sum_{m^{\prime}=m+1}^{k-1}\beta_{k,m,l}^{(1)}\beta_{k,m^{\prime},l^{\prime}}^{(1)}\sum_{j=k+1}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$
$\mathcal{S}_{l,l^{\prime}}^{n,4}:=$ $\sum_{m^{\prime}=l+1}^{n-1}\sum_{m=0}^{m^{\prime}-1}\sum_{k=l\lor m+1}^{m^{\prime}}\beta_{k,m,l}^{(1)}\beta_{k,m^{\prime},l^{\prime}}^{(1)}\sum_{j=m^{\prime}+1}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$
$\mathcal{S}_{l,l^{\prime}}^{n,5}:=$ $\sum_{m=l}^{n-1}\sum_{m^{\prime}=l-l^{\prime}}^{m-1}\sum_{k=l}^{m\land(m^{\prime}+l^{\prime})}\beta_{k,m,l}^{(1)}\beta_{k,m^{\prime},l^{\prime}}^{(1)}\sum_{j=m+1}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$
$\mathcal{S}_{l,l^{\prime}}^{n,6}:=$ $\sum_{m=l}^{n-1}\sum_{k=l}^{m}[\beta_{k,m,l}^{(1)}\beta_{k,m,l^{\prime}}^{(1)}-\mathbb{E}(\beta_{k,m,l}^{(1)}\beta_{k,m,l^{\prime}}^{(1)})]\sum_{j=m+1}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$
$\mathcal{S}_{l,l^{\prime}}^{n,7}:=$ $\sum_{m^{\prime}=l+1}^{n-1}\sum_{m=l}^{m^{\prime}-1}\sum_{k=l}^{m}\beta_{k,m,l}^{(1)}\beta_{k,m^{\prime},l^{\prime}}^{(1)}\sum_{j=m^{\prime}+1}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$
We will only check that $\mathcal{S}_{l,l^{\prime}}^{n,1}$ and $\mathcal{S}_{l,l^{\prime}}^{n,2}$ satisfy (70) since the other terms can be analyzed in the same way. Set $\theta_{k}:=\sum_{j=k+1}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$ and note that by the independent scattered property of $L$ it holds that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,1}\rvert^{2})\leq$ $\sum_{m=0}^{n-1}\sum_{k=l\lor m+1}^{(n-1)\land(m+l)}\sum_{p=l\lor m+1}^{(n-1)\land(m+l)}\mathbb{E}(\beta_{k,m,l}^{(1)}\beta_{k,m,l^{\prime}}^{(1)}\beta_{p,m,l}^{(1)}\beta_{p,m,l^{\prime}}^{(1)})\theta_{k}\theta_{p}$
Lemma 1 implies now that
$\mathbb{E}(\beta_{k,m,l}^{(1)}\beta_{k,m,l^{\prime}}^{(1)}\beta_{p,m,l}^{(1)}\beta_{p,m,l^{\prime}}^{(1)})\lesssim_{a,\mathfrak{t}_{L}}\mathrm{Leb}(\cup_{i=0}^{k\land p-l}\mathcal{P}_{A}^{n}(i,m)).$
This estimate along with the fact that $\rvert\theta_{k}\rvert\leq\int_{t_{k}}^{t_{k+1}}a(s)\mathrm{d}s\leq a(0)\Delta_{n}$ lead to
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,1}\rvert^{2})\lesssim\Delta_{n}\sum_{m=0}^{n-1}\sum_{k=l\lor m+1}^{(n-1)\land(m+l)}\theta_{k}\sum_{p=l\lor m+1}^{k}\mathrm{Leb}(\cup_{i=0}^{p-l}\mathcal{P}_{A}^{n}(i,m)).$
Since
$\sum_{p=l\lor m+1}^{k}\mathrm{Leb}(\cup_{i=0}^{p-l}\mathcal{P}_{A}^{n}(i,m))=\sum_{p=l\lor m+1}^{k}\int_{t_{m}+t_{l}-t_{p}}^{t_{m+1}+t_{l}-t_{p}}a(s)\mathrm{d}s\leq\mathrm{Leb}(A),$
and
$\sum_{k=l\lor m+1}^{(n-1)\land(m+l)}\theta_{k}$ $=\sum_{k=l\lor m+1}^{(n-1)\land(m+l)}\sum_{j=k+1}^{(n-1)\land(m+l)}\mathbb{E}[\alpha_{k+1,j}^{2}]+\sum_{k=l\lor m+1}^{(n-1)\land(m+l)}\sum_{j=(n-1)\land(m+l)}^{n}\mathbb{E}[\alpha_{k+1,j}^{2}]$
$\leq l\Delta_{n}+\mathrm{Leb}(A),$
where in the second inequality we also used (21). This shows that $\mathcal{S}_{l,l^{\prime}}^{n,1}$ satisfies (70). Similar arguments give that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,2}\rvert^{2})=$ $\sum_{m=l-l^{\prime}}^{n-1}\sum_{m^{\prime}=(l-l^{\prime})\lor(m-l^{\prime})+1}^{m-1}\sum_{k,p=l\lor m+1}^{(n-1)\land(m^{\prime}+l^{\prime})}\mathrm{Leb}(\cup_{i=0}^{k\land p-l}\mathcal{P}_{A}^{n}(i,m))\mathrm{Leb}(\cup_{i=0}^{k\land p-l^{\prime}}\mathcal{P}_{A}^{n}(i,m^{\prime}))\theta_{k}\theta_{p}$
$\lesssim$ $\Delta_{n}\sum_{m=l-l^{\prime}}^{n-1}\sum_{m^{\prime}=(l-l^{\prime})\lor(m-l^{\prime})+1}^{m-1}\sum_{k=l\lor m+1}^{(n-1)\land(m^{\prime}+l^{\prime})}\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,m^{\prime}))\theta_{k}$
$\lesssim$ $\Delta_{n}\sum_{m=l-l^{\prime}}^{n-1}\sum_{k=l\lor m+1}^{(n-1)\land(m+l^{\prime})}\theta_{k}\lesssim n\Delta_{n}(l^{\prime}\Delta_{n}+\mathrm{Leb}(A)).$
$\ell=4:$
Recall that $\varsigma_{k,j}:=\sum_{m=k+1}^{j-1}\alpha_{k+1,m}$ and note that as above we may write
$\mathcal{V}_{n}^{(4,4)}$ $=\mathbb{E}(\mathcal{V}_{n}^{(4,4)})+\Delta_{n}^{2}\sum_{l=0}^{n-1}(\overline{c}_{l}^{n})^{2}\mathcal{S}_{l,l}^{n}$
$+2\Delta_{n}^{2}\sum_{l=1}^{n-1}\sum_{l^{\prime}=0}^{l-1}\overline{c}_{l}^{n}\overline{c}_{l^{\prime}}^{n}\frac{\mathcal{S}_{l,l^{\prime}}^{n}}{n\Delta_{n}},$
where
$\mathcal{S}_{l,l^{\prime}}^{n}=$ $\sum_{k=l}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{n-1}\sum_{j=k\lor k^{\prime}+1}^{n}[\varsigma_{k^{\prime},j}\varsigma_{k,j}-\mathbb{E}(\varsigma_{k^{\prime},j}\varsigma_{k,j})]\theta_{k,l,k^{\prime},l^{\prime}}^{j}$
$=$ $\sum_{k=l}^{n-1}\sum_{j=k+1}^{n}[\varsigma_{k,j}^{2}-\mathbb{E}(\varsigma_{k,j})]\theta_{k,l,k^{\prime},l^{\prime}}^{j}$
$+\sum_{k=l+1}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{k-1}\sum_{j=k+1}^{n}\varsigma_{k^{\prime},j}\varsigma_{k,j}\theta_{k,l,k^{\prime},l^{\prime}}^{j}$
$+\sum_{k^{\prime}=l+1}^{n-1}\sum_{k=l}^{k^{\prime}-1}\sum_{j=k^{\prime}+1}^{n}\varsigma_{k^{\prime},j}\varsigma_{k,j}\theta_{k,l,k^{\prime},l^{\prime}}^{j}$
$=:$ $\sum_{\imath=1}^{3}\mathcal{S}_{l,l^{\prime}}^{n,\imath},$
wherein we have let $\theta_{k,l,k^{\prime},l^{\prime}}^{j}:=\mathrm{Leb}(\cup_{i=0}^{k-l\land k^{\prime}-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))$. In this case we are going to show that (68) is satisfied with $\upsilon_{n}=C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)}.$ By construction $\varsigma_{k,j}$ is $\mathscr{F}_{k+1}^{n}$-measurable and independent of $\mathscr{F}_{k}^{n}$, reason why $\mathcal{S}_{l,l^{\prime}}^{n,\imath}$ is a sum of martingale difference for $\imath=1,2,3$. This, along with the fact that $\mathbb{E}(\varsigma_{k,j}^{2})\leq a(0)\Delta_{n}$ and $\sum_{j=k\lor k^{\prime}+1}^{n}\theta_{k,l,k^{\prime},l^{\prime}}^{j}\leq\int_{\rvert t_{k}-t_{k^{\prime}}\rvert}^{\infty}a(s)\mathrm{d}s$ imply that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n}\rvert^{2})\lesssim$ $\Delta_{n}^{2}\sum_{k=1}^{n-1}\sum_{k^{\prime}=0}^{k-1}(\int_{t_{k}-t_{k^{\prime}}}^{\infty}a(s)\mathrm{d}s)^{2}\leq n\Delta_{n}\int_{0}^{n\Delta_{n}}(\int_{u}^{\infty}a(s)\mathrm{d}s)^{2}\mathrm{d}u.$
We now can follow the same arguments as for the case $\ell=2$.
Conditional Covariances
By symmetry, we only need to focus on $\mathcal{V}_{n}^{(\ell,\ell^{\prime})}$ for $\ell=1,2,3,4$ and $\ell^{\prime}\geq\ell+1$.
$\ell=1,\ell^{\prime}=2:$
As it is customary, we set
$\mathcal{S}_{l,l^{\prime}}^{n}$ $=\mathfrak{K}_{3}\sum_{k=l^{\prime}}^{n-1}\chi_{k}\sum_{j=l\lor k+1}^{n-1}\mathrm{Leb}(\cup_{i=0}^{(k-l^{\prime})\land(j-l)}\mathcal{P}_{A}^{n}(i,j))+\sum_{j=l\lor l^{\prime}}^{n-1}\sum_{k=l^{\prime}}^{j-1}[\beta_{j,j-1,l}^{(2)}\chi_{k}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\chi_{k})]\mathbb{E}[(\beta_{k,j,l^{\prime}}^{(1)})^{2}]$
$=:\mathcal{S}_{l,l^{\prime}}^{n,1}+\mathcal{S}_{l,l^{\prime}}^{n,2}$
This case will be slightly different from the others. First observe that in view of
$\sum_{j=l\lor k+1}^{n-1}\mathrm{Leb}(\cup_{i=0}^{(k-l^{\prime})\land(j-l)}\mathcal{P}_{A}^{n}(i,j))\leq\mathrm{Leb}(A),$
we conclude as in previous cases that $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,1}\rvert^{2})\lesssim n\Delta_{n}$. Second, decompose
$\mathcal{S}_{l,l^{\prime}}^{n,2}=\mathcal{S}_{l,l^{\prime}}^{\star,n}+\tilde{\mathcal{S}}_{l,l^{\prime}}^{n,2},$
where
$\mathcal{S}_{l,l^{\prime}}^{\star,n}:=\sum_{j=l\lor l^{\prime}}^{l+l^{\prime}}\sum_{k=l^{\prime}}^{j-1}[\beta_{j,j-1,l}^{(2)}\chi_{k}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\chi_{k})]\mathbb{E}[(\beta_{k,j,l^{\prime}}^{(1)})^{2}].$
By arguing exactly in the same way of the proof of Lemma 3 in [17] (specifically the analysis done for $\frac{1}{n\Delta_{n}}\sum_{j=l_{n}}^{n}\mathbb{E}(\xi_{j,n}^{(2)}\xi_{j,n}^{(3)}\mid\mathscr{G}_{j-1}^{n})$ under the notation introduced in that paper) we conclude that $\tilde{\mathcal{S}}_{l,l^{\prime}}^{n,2}$ satisfy (69) and that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{\star,n}\rvert)\lesssim\Delta_{n}(l\lor l^{\prime}).$ (72)
Note that by replacing (69) by (72) in the arguments above, we obtain that $\mathbb{E}(\rvert\mathcal{V}_{n}^{(1,2)}-\mathbb{E}(\mathcal{V}_{n}^{(1,2)})\rvert)\rightarrow 0$, as $n\rightarrow\infty$.
$\ell=1,\ell^{\prime}=3:$
In this case, by applying (71), we deduce that (67) holds with
$\mathcal{S}_{l,l^{\prime}}^{n}$ $=-\sum_{m=0}^{n-1}\sum_{k=l^{\prime}}^{(n-l-1)\land(m+l^{\prime})}\beta_{k,m,l^{\prime}}^{(1)}\theta_{m,k}^{l,l^{\prime}}-\sum_{j=l\lor l^{\prime}+1}^{n-1}\sum_{k=l^{\prime}}^{j-1}\beta_{j,j-1,l}^{(2)}\beta_{k,j-1,l^{\prime}}^{(2)}\mathbb{E}(\alpha_{k+1,j}^{2})$
$=:\mathcal{S}_{l,l^{\prime}}^{n,1}+\mathcal{S}_{l,l^{\prime}}^{n,2}$
where $\theta_{m,k}^{l,l^{\prime}}:=\sum_{j=m\lor(k+l)+1}^{n-1}\mathbb{E}(\alpha_{k+1,j}^{3})$. By arguing exactly in the same way as $\ell=\ell^{\prime}=3$, we deduce that $\mathcal{S}_{l,l^{\prime}}^{n,1}$ satisfies the bound in (69). Now, assume without loss of generality that $l\geq l^{\prime}$ and put
$\vartheta_{j,l,l^{\prime}}^{(1)}$ $=\sum_{m=0}^{j-1-l}\sum_{i=0}^{m}\alpha_{i,m}\sum_{k=i+l^{\prime}}^{m+l^{\prime}}\mathbb{E}(\alpha_{k+1,j}^{2});$
$\vartheta_{j,l,l^{\prime}}^{(2)}$ $=\sum_{m=j-l}^{j-1-l^{\prime}}\sum_{i=0}^{m}\alpha_{i,m}\sum_{k=i+l^{\prime}}^{m+l^{\prime}}\mathbb{E}(\alpha_{k+1,j}^{2});$
$\vartheta_{j,l,l^{\prime}}^{(3)}$ $=\sum_{i=0}^{j-1-l^{\prime}}\sum_{m=j-l^{\prime}}^{j-1}\alpha_{i,m}\sum_{k=i+l^{\prime}}^{j-1}\mathbb{E}(\alpha_{k+1,j}^{2}).$
Then, we can further decompose $\mathcal{S}_{l,l^{\prime}}^{n,2}=\sum_{\imath=1}^{3}\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,\imath}$, in which $\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,1}:=-\sum_{j=l\lor l^{\prime}+1}^{n-1}\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(1)}$ and
$\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,\imath}:=-\sum_{j=l\lor l^{\prime}+1}^{n-1}[\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(\imath)}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(\imath)})],\,\,\,\imath=2,3.$
Note that $\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(\imath)}$ are $l$-dependent. Furthermore, by Rosenthal’s inequality, Lemma 1, and the fact that for $m\leq j-1-l^{\prime}$ it holds that $\sum_{k=i+l^{\prime}}^{m+l^{\prime}}\mathbb{E}(\alpha_{k+1,j}^{2})\lesssim\int_{t_{j}-t_{m+l^{\prime}}}^{t_{j+1}-t_{m+l^{\prime}}}a(s)\mathrm{d}s,$ we infer the following estimate
$\mathbb{E}(\rvert\vartheta_{j,l,l^{\prime}}^{(\imath)}\rvert^{4})\lesssim\Delta_{n}^{4},\,\,\imath=1,2.$
Likewise, $\mathbb{E}(\rvert\vartheta_{j,l,l^{\prime}}^{(3)}\rvert^{4})\lesssim\Delta_{n}^{4}$. All in all, allow us to conclude as in the case $\ell=\ell^{\prime}=1$ that $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,2}\rvert^{2})\lesssim n\Delta_{n}(l\Delta_{n})$.
$\ell=1,\ell^{\prime}=4:$
Here we will argue as in the situation when $\ell=1,\ell^{\prime}=2$. We start by noting that (67) is satisfied with
$\mathcal{S}_{l,l^{\prime}}^{n}$ $=-\mathfrak{K}_{3}\sum_{k=l^{\prime}}^{n-1}\sum_{j=l\lor k+1}^{n-1}\varsigma_{k,j}\mathrm{Leb}(\cup_{i=0}^{(k-l^{\prime})\land(j-l)}\mathcal{P}_{A}^{n}(i,j))$
$-\sum_{j=l\lor l^{\prime}}^{n-1}\sum_{k=l^{\prime}}^{j-1}[\beta_{j,j-1,l}^{(2)}\varsigma_{k,j}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\varsigma_{k,j})]\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))$
$=:\mathcal{S}_{l,l^{\prime}}^{n,1}+\mathcal{S}_{l,l^{\prime}}^{n,2}.$
Furthermore, by using that $\sum_{j=l\lor k+1}^{n-1}\mathrm{Leb}(\cup_{i=0}^{(k-l^{\prime})\land(j-l)}\mathcal{P}_{A}^{n}(i,j))\leq\mathrm{Leb}(A)$ and $\mathbb{E}(\varsigma_{k,p}\varsigma_{k,j})\lesssim\Delta_{n}$, we conclude as in the case $\ell=\ell^{\prime}=4$ that $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,1}\rvert^{2})\lesssim n\Delta_{n}$. Next, we decompose
$\mathcal{S}_{l,l^{\prime}}^{n,2}=\mathcal{S}_{l,l^{\prime}}^{\star,n}+\tilde{\mathcal{S}}_{l,l^{\prime}}^{n,2},$
where
$\mathcal{S}_{l,l^{\prime}}^{\star,n}:=\sum_{j=l\lor l^{\prime}}^{l+l^{\prime}+1}\sum_{k=l^{\prime}}^{j-1}[\beta_{j,j-1,l}^{(2)}\varsigma_{k,j}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\varsigma_{k,j})]\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j)).$
The Cauchy-Swarchz inequality and the fact that $\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))=\int_{t_{j}-t_{k-l^{\prime}}}^{t_{j+1}-t_{k-l^{\prime}}}a(s)\mathrm{d}s$ imply that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{\star,n}\rvert)\lesssim\Delta_{n}\sum_{j=l\lor l^{\prime}}^{l+l^{\prime}+1}\left(\sum_{k=l^{\prime}}^{j-1}\int_{t_{j}-t_{k-l^{\prime}}}^{t_{j+1}-t_{k-l^{\prime}}}a(s)\mathrm{d}s\right)^{1/2}\lesssim\Delta_{n}(l\lor l^{\prime}).$
On the other hand, by setting
$\vartheta_{j,l,l^{\prime}}^{(1)}$ $=\sum_{m=l^{\prime}+1}^{j-l-1}\sum_{k=l^{\prime}}^{m-1}\alpha_{k+1,m}\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j));$
$\vartheta_{j,l,l^{\prime}}^{(2)}$ $=\sum_{m=j-l}^{j-1}\sum_{k=l^{\prime}}^{m-1}\alpha_{k+1,m}\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j)).$
we can further decompose $\tilde{\mathcal{S}}_{l,l^{\prime}}^{n,2}=\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,1}+\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,2}$, where $\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,1}:=-\sum_{j=l\lor l^{\prime}+1}^{n-1}\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(1)}$ and
$\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,2}:=-\sum_{j=l\lor l^{\prime}+1}^{n-1}[\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(2)}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(2)})].$
As in the previous case, $\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(\imath)}$ are $l$-dependent and $\mathbb{E}(\rvert\vartheta_{j,l,l^{\prime}}^{(\imath)}\rvert^{4})\lesssim\Delta_{n}^{4}$, reason why $\mathbb{E}(\rvert\tilde{\mathcal{S}}_{l,l^{\prime}}^{n,2}\rvert^{2})\lesssim n\Delta_{n}(l\Delta_{n})$.
$\ell=2,\ell^{\prime}=3:$
Another application of (71) gives us that in this situation that
$\mathcal{S}_{l,l^{\prime}}^{n}=$ $\sum_{k=l^{\prime}+l+1}^{n-1}\chi_{k}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}\beta_{k^{\prime},k-1,l^{\prime}}^{(2)}\theta_{k,k^{\prime}}+\sum_{k=l^{\prime}+l+1}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}[\chi_{k}\beta_{k^{\prime},k,l^{\prime}}^{(1)}-\mathbb{E}(\chi_{k}\beta_{k^{\prime},k,l^{\prime}}^{(1)})]\theta_{k,k^{\prime}}$
$+\sum_{m=l^{\prime}+l+2}^{n-2}\sum_{k=l^{\prime}+l+1}^{m-1}\chi_{k}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}\beta_{k^{\prime},m,l^{\prime}}^{(1)}\theta_{m,k^{\prime}}$
$=:$ $\sum_{\imath=1}^{3}\mathcal{S}_{l,l^{\prime}}^{n,\imath},$
where $\theta_{k,k^{\prime}}:=\sum_{j=k+1}^{n-1}\mathbb{E}[\alpha_{k^{\prime}+1,j}^{2}]$. Once again, each $\mathcal{S}_{l,l^{\prime}}^{n,\imath}$ is a sum of martingale difference. Furthermore, by reasoning as in the proof of Lemma 3 in [17] (see the discussion concerned to $\frac{1}{n\Delta_{n}}\sum_{j=l_{n}}^{n}\mathbb{E}(\xi_{j,n}^{(3)}\xi_{j,n}^{(4)}\mid\mathscr{G}_{j-1}^{n})$), the second moment of each summand in $\mathcal{S}_{l,l^{\prime}}^{n,\imath}$ is bounded up to a constant by $\Delta_{n}$, from which it is clear that $\mathcal{S}_{l,l^{\prime}}^{n}$ satisfies (68).
$\ell=2,\ell^{\prime}=4:$
Obviously in this situation we have
$\mathcal{S}_{l,l^{\prime}}^{n}=-\sum_{j=l\lor l^{\prime}}^{n-1}\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}[\chi_{k}\varsigma_{k^{\prime},j}-\mathbb{E}(\chi_{k}\varsigma_{k^{\prime},j})]\mathrm{Leb}(\cup_{i=0}^{(k^{\prime}-l^{\prime})\land(k-l)}\mathcal{P}_{A}^{n}(i,j)).$
Assume now that $l\geq l^{\prime}$. Then, using that for $k\geq k^{\prime}+1$, it holds
$\chi_{k}=L(\cup_{i=0}^{k^{\prime}}\mathcal{P}_{A}^{n}(i,k))+\sum_{i=k^{\prime}+1}^{k}\alpha_{i,k}=:\lambda_{k^{\prime},k}+\sum_{i=k^{\prime}+1}^{k}\alpha_{i,k}$
we can write $\mathcal{S}_{l,l^{\prime}}^{n}=\sum_{\imath=1}^{4}\mathcal{S}_{l,l^{\prime}}^{n,\imath},$ where we have set
$\mathcal{S}_{l,l^{\prime}}^{n,1}$ $:=-\sum_{k^{\prime}=l}^{n-2}\sum_{m=k^{\prime}+1}^{n-2}\alpha_{k^{\prime}+1,m}\sum_{k=l}^{k^{\prime}}\chi_{k}\bar{\theta}_{m,k^{\prime},l^{\prime}};$
$\mathcal{S}_{l,l^{\prime}}^{n,2}$ $:=-\sum_{j=l}^{n-1}\sum_{k^{\prime}=l}^{j-2}\sum_{i=k^{\prime}+1}^{j-1}\sum_{k=i}^{j-1}[\varsigma_{k^{\prime},j}\alpha_{i,k}-\mathbb{E}(\varsigma_{k^{\prime},j}\alpha_{i,k})]\theta_{k,l,k^{\prime},l^{\prime}}^{j};$
$\mathcal{S}_{l,l^{\prime}}^{n,3}$ $:=-\sum_{k^{\prime}=l}^{n-3}\sum_{m=k^{\prime}+1}^{n-2}\alpha_{k^{\prime}+1,m}\sum_{k=k^{\prime}+1}^{n-2}\lambda_{k^{\prime},k}\bar{\theta}_{m\lor k,k^{\prime},l^{\prime}};$
$\mathcal{S}_{l,l^{\prime}}^{n,4}$ $:=-\sum_{j=l}^{n-1}\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{l-1}[\chi_{k}\varsigma_{k^{\prime},j}-\mathbb{E}(\chi_{k}\varsigma_{k^{\prime},j})]\theta_{k,l,k^{\prime},l^{\prime}}^{j},$
in which $\theta_{k,l,k^{\prime},l^{\prime}}^{j}:=\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k^{\prime}-l^{\prime})}\mathcal{P}_{A}^{n}(i,j))$ and $\bar{\theta}_{m,k^{\prime},l^{\prime}}:=\mathrm{Leb}(\cup_{i=0}^{(k^{\prime}-l^{\prime})\land(k-l)}\cup_{j=m+1}^{n-1}\mathcal{P}_{A}^{n}(i,j))\leq\mathrm{Leb}(A)$. The same line of arguments used in the case $\ell=\ell^{\prime}=2$ and $\ell=\ell^{\prime}=4$ allow us to conclude that $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\imath}\rvert^{2})\lesssim n\Delta_{n}(C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)})$, for $\imath=1,2$, respectively. Next, note that $\mathcal{S}_{l,l^{\prime}}^{n,3}$ is once again a sum of martingale differences. Furthermore, for a given fix $k^{\prime}$, $(\alpha_{k^{\prime}+1,m})_{k^{\prime}+1\leq m\leq n-2}$ and $(\lambda_{k^{\prime},k})_{k^{\prime}+1\leq k\leq n-2}$ are independent due to the independent scattered property of $L$. Therefore
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,3}\rvert^{2})=$ $\sum_{k^{\prime}=l}^{n-3}\sum_{m=k^{\prime}+1}^{n-2}\mathbb{E}(\alpha_{k^{\prime}+1,m}^{2})\sum_{k=k^{\prime}+1}^{n-2}\mathbb{E}(\lambda_{k^{\prime},k}^{2})(\bar{\theta}_{m\lor k,k^{\prime},l^{\prime}})^{2}$
$\lesssim$ $\sum_{k^{\prime}=l}^{n-3}\sum_{m=k^{\prime}+1}^{n-2}\mathbb{E}(\alpha_{k^{\prime}+1,m}^{2})\lesssim n\Delta_{n}.$
After some tedious, yet simple, change of order of summations we get that
$\mathcal{S}_{l,l^{\prime}}^{n,4}=$ $-\sum_{k^{\prime}=l^{\prime}}^{l-1}\sum_{k=l}^{n-2}[\alpha_{k^{\prime}+1,k}^{2}-\mathbb{E}(\alpha_{k^{\prime}+1,k}^{2})]\bar{\theta}_{k,k^{\prime},l^{\prime}}$
$-\sum_{k^{\prime}=l^{\prime}}^{l-1}\sum_{k=l}^{n-2}\sum_{m=k^{\prime}+1}^{n-2}\alpha_{k^{\prime}+1,k}\alpha_{k^{\prime}+1,m}\bar{\theta}_{m\lor k,k^{\prime},l^{\prime}}\mathbf{1}_{m\geq k+1}$
$-\sum_{k^{\prime}=l^{\prime}}^{l-1}\sum_{k=l}^{n-2}\sum_{m=k^{\prime}+1}^{n-2}\alpha_{k^{\prime}+1,k}\alpha_{k^{\prime}+1,m}\bar{\theta}_{m\lor k,k^{\prime},l^{\prime}}\mathbf{1}_{m\leq k-1}$
$-\sum_{k^{\prime}=l^{\prime}}^{l-1}\sum_{m=k^{\prime}+1}^{n-2}\alpha_{k^{\prime}+1,m}\sum_{i=0}^{k^{\prime}}\sum_{k=l}^{n-2}\alpha_{i,k}\bar{\theta}_{m\lor k,k^{\prime},l^{\prime}}$
$-\sum_{k^{\prime}=l^{\prime}}^{l-1}\sum_{m=k^{\prime}+1}^{n-2}\alpha_{k^{\prime}+1,m}\sum_{k=l}^{n-2}\sum_{i=k^{\prime}+2}^{k}\alpha_{i,k}\bar{\theta}_{m\lor k,k^{\prime},l^{\prime}}.$
Using this we conclude that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,4}\rvert^{2})\lesssim$ $\mathrm{Leb}(A)+\sum_{k^{\prime}=l^{\prime}}^{l-1}\sum_{k=l}^{n-2}\sum_{m=k^{\prime}+1}^{k}\mathbb{E}(\alpha_{k^{\prime}+1,k}^{2})\mathbb{E}(\alpha_{k^{\prime}+1,m}^{2})$
$+\sum_{k^{\prime}=l^{\prime}}^{l-1}\sum_{m=k^{\prime}+1}^{n-2}\mathbb{E}(\alpha_{k^{\prime}+1,m}^{2})\sum_{i=0}^{k^{\prime}}\sum_{k=l}^{n-2}\mathbb{E}(\alpha_{i,k}^{2})$
$\lesssim$ $\mathrm{Leb}(A)(1+\Delta_{n}+l\Delta_{n}).$
Analogous arguments can be used for the case $l<l^{\prime}$.
$\ell=3,\ell^{\prime}=4:$
In this case it is enough to note that by swapping $k$ with $k^{\prime}$ and replacing $\chi_{k}$ by $\varsigma_{k^{\prime},j}$ in the case $\ell=2,\ell^{\prime}=3$, we deduce that (67) holds and $\mathcal{S}_{l,l^{\prime}}^{n}$ satisfies (68).
In the proof of the next lemma we will often use the following local approximation of $a$:
$a_{n}(s):=\int_{0}^{1}a(\Delta_{n}x+s)\mathrm{d}x.$ (73)
Clearly, $a_{n}\leq a$ and $a_{n}\rightarrow a$ uniformly on compacts, thanks to Assumption 1. We recall the reader that $\tau_{n}(s):=[s/\Delta_{n}]\Delta_{n}$, $s\geq 0$.
Lemma 8.
Let Assumptions of Theorems 5 and 6 hold. Then,
$\Sigma_{n}^{(\ell,\ell^{\prime})}(l,l^{\prime}):=\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}}^{n}\mathbb{E}(\xi_{j,l,n}^{(\ell)}\xi_{j,l^{\prime},n}^{(\ell^{\prime})}\mid\mathscr{G}_{j-1}^{n})=F_{n}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})+\mathrm{o}(1),\,\,\ell,\ell^{\prime}=1,2,3,4,$ (74)
uniformly on $l,l^{\prime}$ and where $F_{n}^{(\ell,\ell^{\prime})}$ is a sequence of bounded functions such that for all $s,r\geq 0$, $F_{n}^{(\ell,\ell^{\prime})}(\tau_{n}(s),\tau_{n}(r))\rightarrow F^{(\ell,\ell^{\prime})}(s,r)$ as $n\rightarrow\infty$ for some measurable function $F^{(\ell,\ell^{\prime})}$ satisfying that $\sum_{\ell,\ell^{\prime}=1}^{4}F_{n}^{(\ell,\ell^{\prime})}(\tau_{n}(s),\tau_{n}(r))\rightarrow\sum_{\ell,\ell^{\prime}=1}^{4}F^{(\ell,\ell^{\prime})}(s,r)=\Sigma_{a}(s,r),$for all $s,r$.
Proof.
As in the proof of Lemma 7 we analyze separately the terms $\Sigma_{n}^{(\ell,\ell^{\prime})}(l,l^{\prime})$. Note that in view that $\Sigma_{n}^{(\ell^{\prime},\ell)}(l,l^{\prime})=\Sigma_{n}^{(\ell,\ell^{\prime})}(l^{\prime},l)$, it is enough to show that (74) holds for $\ell=1,2,3,4,$ and $\ell^{\prime}\geq\ell$. Furthermore, it will be implicitly assumed that for fixed $n$, the function $F_{n}^{(\ell,\ell^{\prime})}$ defined below vanishes outside of $0,n\Delta_{n}]$. Finally, unless said otherwise, the error terms (represented as usual by $\mathrm{O}(\cdot)$) presented below will always be uniformly on $l,l^{\prime}$.
$\ell=\ell^{\prime}=1:$
Using (63) and Lemma 1, we get that
$\Sigma_{n}^{(1,1)}(l,l^{\prime})=$ $\frac{\mathfrak{K}_{4}}{\Delta_{n}n}\sum_{j=l\lor l^{\prime}}^{n-1}\mathrm{Leb}(\cup_{i=0}^{j-l\lor l^{\prime}}\mathcal{P}_{A}^{n}(i,j))$
$+a_{n}(0)\frac{1}{n}\sum_{j=l\lor l^{\prime}}^{n-1}\mathrm{Leb}(\cup_{i=0}^{j-l\lor l^{\prime}}\cup_{m=j-l\land l^{\prime}}^{j-1}\mathcal{P}_{A}^{n}(i,m))$
$-\frac{1}{\Delta_{n}n}\sum_{j=l\lor l^{\prime}}^{n-1}\mathrm{Leb}(\cup_{i=0}^{j-l}\mathcal{P}_{A}^{n}(i,j))\mathrm{Leb}(\cup_{i=0}^{j-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))+\mathrm{O}(\Delta_{n})$
$=$ $F_{n,1}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})+F_{n,2}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})+\mathrm{O}(\Delta_{n}).$
where for $s,r\leq n\Delta_{n}$
$F_{n,1}^{(1,1)}(s,r)$ $=\mathfrak{K}_{4}\frac{1}{\Delta_{n}}\int_{s\lor r}^{s\lor r+\Delta_{n}}a(u)\mathrm{d}u\left(1-\frac{s\lor r}{n\Delta_{n}}\right)$
$=\mathfrak{K}_{4}a_{n}(s\lor r)\left(1-\frac{s\lor r}{n\Delta_{n}}\right)$
and
$F_{n,2}^{(1,1)}(s,r)=a_{n}(0)\int_{s\lor r-r\land s}^{s\lor r}a(u)\mathrm{d}u\left(1-\frac{s\lor r}{n\Delta_{n}}\right).$
Thus, in view that $\frac{\tau_{n}(r)\lor\tau_{n}(s)}{n\Delta_{n}}\leq 1$ for $r,s\leq n\Delta_{n}$ and $a_{n}\leq a\leq a(0)$, we easily conclude that $\rvert F_{n,1}^{(1,1)}(s,r)\rvert\leq 2\mathfrak{K}_{4}a(0)$ as well as $\rvert F_{n,2}^{(1,1)}(s,r)\rvert\leq 2a(0)\mathrm{Leb}(A).$ This shows that the function $F_{n}^{(1,1)}:=F_{n,1}^{(1,1)}+F_{n,2}^{(1,1)}$ is bounded and for every given $s,r\geq 0$
$F_{n}^{(1,1)}(\tau_{n}(s),\tau_{n}(r))$ $\rightarrow\mathfrak{K}_{4}a(s\lor r)+a(0)\int_{s\lor r-s\land r}^{s\lor r}a(u)\mathrm{d}u=:F^{(1,1)}(s,r),$
just as required.
$\ell=\ell^{\prime}=2:$
Just as above we get that
$\Sigma_{n}^{(2,2)}(l,l^{\prime})=$ $a_{n}(0)\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{k=l\lor l^{\prime}}^{j-1}\mathrm{Leb}(\cup_{i=0}^{k-l\lor l^{\prime}}\mathcal{\mathcal{P}}_{A}^{n}(i,j))$
$=$ $a_{n}(0)\frac{1}{n}\sum_{j=l\lor l^{\prime}+1}^{n}\int_{t_{l+1}\lor t_{l^{\prime}+1}}^{t_{j+1}}a(u)\mathrm{d}u.$
Thus, (74) holds in this situation with
$F_{n}^{(2,2)}(s,r)$ $=a_{n}(0)\frac{1}{n}\sum_{j=[s/\Delta_{n}]\lor[r/\Delta_{n}]+1}^{n}\int_{s\lor r}^{t_{j}}a(u)\mathrm{d}u.$ (75)
Clearly $\rvert F_{n}^{(2,2)}(s,r)\rvert\leq a(0)\mathrm{Leb}(A)$ and by a simple change of variables and the DCT we further have that $F_{n}^{(2,2)}(\tau_{n}(s),\tau_{n}(r))\rightarrow a(0)\int_{s\lor r}^{\infty}a(u)\mathrm{d}u=:F^{(2,2)}(s,r),$ as needed.
$\ell=\ell^{\prime}=3:$
In this situation we have due to (63) that
$\Sigma_{n}^{(3,3)}(l,l^{\prime})=\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{k=l\lor l^{\prime}}^{j-1}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k+1,j))\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{j}-t_{k}+t_{l}\lor t_{l^{\prime}}}a(u)\mathrm{d}u,$
By letting $m=j-k+l\lor l^{\prime}$ and denoting $A_{m}^{l,l^{\prime}}:=\int_{t_{m-1}-t_{l}\lor t_{l^{\prime}}}^{t_{m}-t_{l}\lor t_{l^{\prime}}}a(s)\mathrm{d}s$, we deduce that
$\sum_{k=l\lor l^{\prime}}^{j-1}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k+1,j))\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{j}-t_{k}+t_{l}\lor t_{l^{\prime}}}a(s)\mathrm{d}s$ $=\sum_{m=l\lor l^{\prime}+1}^{j}\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{m}}a(u)\mathrm{d}u[A_{m}^{l,l^{\prime}}-A_{m+1}^{l,l^{\prime}}]$
$=\Delta_{n}a_{n}(0)\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{l+1}\lor t_{l^{\prime}+1}}a(u)\mathrm{d}u$
$+\sum_{m=l\lor l^{\prime}+2}^{j}\int_{t_{m-1}}^{t_{m}}a(u)\mathrm{d}u\int_{t_{m-1}}^{t_{m}}a(u-t_{l}\lor t_{l^{\prime}})\mathrm{d}u$
$-\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{j}}a(u)\mathrm{d}u\int_{t_{j}}^{t_{j+1}}a(u-t_{l}\lor t_{l^{\prime}})\mathrm{d}u$
$=\Delta_{n}a_{n}(0)\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{l+1}\lor t_{l^{\prime}+1}}a(u)\mathrm{d}u$
$+\Delta_{n}\int_{t_{l+1}\lor t_{l^{\prime}+1}}^{t_{j}}a_{n}(\tau_{n}(u))a(u-t_{l}\lor t_{l^{\prime}})\mathrm{d}u$
$+\mathrm{O}\left(\int_{t_{j}}^{t_{j+1}}a(u-t_{l}\lor t_{l^{\prime}})\mathrm{d}u\right)$
In consequence, we have that $\Sigma_{n}^{(3,3)}(l,l^{\prime})=F_{n}^{(3,3)}(t_{l},t_{l^{\prime}})+\mathrm{O}(1/n\Delta_{n}),$ where
$F_{n}^{(3,3)}(s,r):=\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s\lor r}^{y}a_{n}(\tau_{n}(u))a(u-s\lor r)\mathrm{d}u\mathrm{d}y+a(0)\int_{s\lor r-s\land r}^{s\lor r}a(u)\mathrm{d}u\left(1-\frac{s\lor r}{n\Delta_{n}}\right).$
Invoking again the inequality $a_{n}\leq a\leq a(0)$, we obtain that $\rvert F_{n}^{(3)}(s,r)\rvert\leq 3a(0)\mathrm{Leb}(A)$ and just as for $F_{n}^{(2,2)}$ we also get
$F_{n}^{(3,3)}(\tau_{n}(s),\tau_{n}(r))\rightarrow\int_{s\lor r}^{\infty}a(u)a(u-s\lor r)\mathrm{d}u+a(0)\int_{s\lor r-r\land s}^{s\lor r}a(u)\mathrm{d}u=:F^{(3,3)}(s,r).$
$\ell=\ell^{\prime}=4:$
Repeating the same reasoning as above we get in this case that
$\Sigma_{n}^{(4,4)}(l,l^{\prime})$ $=\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{k=l\lor l^{\prime}}^{j-1}\mathrm{Leb}(\cup_{m=k+1}^{j}\mathcal{\mathcal{P}}_{A}^{n}(k+1,m))\mathrm{Leb}(\cup_{i=0}^{k-l\lor l^{\prime}}\mathcal{P}_{A}^{n}(i,j))$
$=\frac{1}{n}\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{k=l\lor l^{\prime}}^{j-1}\int_{t_{j}-t_{k}}^{t_{j+1}-t_{k}}[a_{n}(0)-a(u)]a_{n}(t_{j}-t_{k}+t_{l}\lor t_{l^{\prime}})\mathrm{d}u$
$=\frac{1}{n}\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{m=l\lor l^{\prime}+1}^{j-1}\int_{t_{m}-t_{l}\lor t_{l^{\prime}}}^{t_{m+1}-t_{l}\lor t_{l^{\prime}}}[a_{n}(0)-a(u)]a_{n}(t_{m})\mathrm{d}u$
$=F_{n}^{(4,4)}(t_{l},t_{l^{\prime}})+\mathrm{O}(\Delta_{n})$
with
$F_{n}^{(4,4)}(s,r):=\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s\lor r}^{y}[a_{n}(0)-a(u-s\lor r)]a_{n}(\tau_{n}(u))\mathrm{d}u\mathrm{d}y,\,\,0\leq r,s\leq n\Delta_{n}.$
It is not difficult to see that $F_{n}^{(4,4)}$ is bounded by $2a(0)(\Delta_{n}a(0)+\mathrm{Leb}(A))$ and arguing as before we also have that
$F_{n}^{(4,4)}(\tau_{n}(s),\tau_{n}(r))\rightarrow\int_{s\lor r}^{\infty}[a(0)-a(u-s\lor r)]a(u)\mathrm{d}u=:F^{(4,4)}(s,r).$
$\ell=1,\ell^{\prime}=2:$
In this case we use relation (64) to get that
$\Sigma_{n}^{(1,2)}(l,l^{\prime})=\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}+1}^{n-1}\sum_{k=l^{\prime}}^{j-1}\mathbb{E}(\chi_{k}\beta_{j,j-1,l}^{(2)})\mathbb{E}[(\beta_{k,j,l^{\prime}}^{(1)})^{2}].$
But
$\sum_{k=l^{\prime}}^{j-1}\mathbb{E}(\chi_{k}\beta_{j,j-1,l}^{(2)})\mathbb{E}[(\beta_{k,j,l^{\prime}}^{(1)})^{2}]$ $=\sum_{k=l^{\prime}}^{j-1}\mathbb{E}[(\beta_{j,k,l}^{(1)})^{2}]\mathbb{E}[(\beta_{k,j,l^{\prime}}^{(1)})^{2}]\mathbf{1}_{k\geq j-l}$
$=\Delta_{n}\sum_{k=l^{\prime}\lor(j-l)}^{j-1}\int_{t_{j}-t_{k}+t_{l^{\prime}}}^{t_{j+1}-t_{k}+t_{l^{\prime}}}a(u)\mathrm{d}ua_{n}(t_{k}-t_{j}+t_{l})$
$=\Delta_{n}\sum_{m=l^{\prime}+1}^{j-(j-l-l^{\prime})^{+}}\int_{t_{m}}^{t_{m+1}}a(u)\mathrm{d}ua_{n}(t_{l}+t_{l^{\prime}}-t_{m}).$
Hence, $\Sigma_{n}^{(1,2)}(l,l^{\prime})=F_{n}^{(1,2)}(t_{l},t_{l^{\prime}})+\mathrm{O}(1/n),$ in which
$F_{n}^{(1,2)}(s,r):=\frac{1}{n}\sum_{j=[s/\Delta_{n}]\lor[r/\Delta_{n}]}^{n}\sum_{m=[r/\Delta_{n}]+1}^{j-(j-[s/\Delta_{n}]-[r/\Delta_{n}])^{+}}\int_{t_{m}}^{t_{m+1}}a(u)a_{n}(s+r-\tau_{n}(u))\mathrm{d}u$
which is obviously bounded by $a(0)\mathrm{Leb}(A)$ and for $n$ large enough and fix $s,r\geq 0$ we further have
$F_{n}^{(1,2)}(s,r)=\int_{r}^{s+r}a(u)a_{n}(s+r-\tau_{n}(u))\mathrm{d}u\left(1-\frac{\tau_{n}(r)+\tau_{n}(s)}{n\Delta_{n}}\right)+\mathrm{O}(\frac{s\lor r}{n\Delta_{n}}).$
Consequently, as $n\rightarrow\infty$
$F_{n}^{(1,2)}(\tau_{n}(s),\tau_{n}(r))\rightarrow\int_{r}^{s+r}a(u)a(s+r-u)\mathrm{d}u=:F^{(1,2)}(s,r),$
by the DCT. This is exactly what we were seeking for.
$\ell=1,\ell^{\prime}=3:$
Exactly as above
$\Sigma_{n}^{(1,3)}(l,l^{\prime})=\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}+1}^{n-1}\sum_{k=l^{\prime}}^{j-1}\mathbb{E}(\beta_{j,j-1,l}^{(2)}\beta_{k,j-1,l^{\prime}}^{(2)})\mathbb{E}(\alpha_{k+1,j}\chi_{j}).$
Using that
$\mathbb{E}(\beta_{j,j-1,l}^{(2)}\beta_{k,j-1,l^{\prime}}^{(2)})=\begin{cases}\sum_{i=0}^{k-l^{\prime}}\mathrm{Leb}(\cup_{m=j-l}^{j-1}\mathcal{\mathcal{P}}_{A}^{n}(i,m)),&\text{if }k\leq j-(l-l^{\prime});\\ \sum_{m=k-l^{\prime}}^{j-1}\mathrm{Leb}(\cup_{i=0}^{j-l}\mathcal{\mathcal{P}}_{A}^{n}(i,m)),&\text{otherwise},\end{cases}$
we decompose
$\sum_{k=l^{\prime}}^{j-1}\mathbb{E}(\beta_{j,j-1,l}^{(2)}\beta_{k,j-1,l^{\prime}}^{(2)})\mathbb{E}(\alpha_{k+1,j}\chi_{j})=\mathfrak{S}_{j,l,l^{\prime}}^{n,1}+\mathfrak{S}_{j,l,l^{\prime}}^{n,2},$
where
$\mathfrak{S}_{j,l,l^{\prime}}^{n,1}$ $:=\sum_{i=0}^{j-(l^{\prime}+1)\lor l}\mathrm{Leb}(\cup_{m=j-l}^{j-1}\mathcal{\mathcal{P}}_{A}^{n}(i,m))\sum_{k=i+l^{\prime}}^{j-1-(l-l^{\prime}+1)^{+}}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k+1,j)),$
$\mathfrak{S}_{j,l,l^{\prime}}^{n,2}$ $:=\sum_{m=j-l+1}^{j-1}\mathrm{Leb}(\cup_{i=0}^{j-l}\mathcal{\mathcal{P}}_{A}^{n}(i,m))\sum_{k=l^{\prime}}^{j-1}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k+1,j))\mathbf{1}_{m+l^{\prime}\geq k\geq j-(l-l^{\prime})+1}.$
Simple algebraic manipulations result in
$\mathfrak{S}_{j,l,l^{\prime}}^{n,1}=$ $\Delta_{n}\sum_{m=l\lor(l^{\prime}+1)}^{j-1}\int_{t_{m}}^{t_{m+1}}[a(u-t_{l})-a(u)][a_{n}((t_{l}-t_{l^{\prime}+1})^{+})-a_{n}(t_{m}-t_{l^{\prime}})]\mathrm{d}u$
$+\mathrm{O}\left(\Delta_{n}\int_{t_{j}-t_{l}}^{\infty}a(u)\mathrm{d}u\right)$
and
$\mathfrak{S}_{j,l,l^{\prime}}^{n,2}$ $=\Delta_{n}\sum_{m=1}^{l-1}\int_{t_{m}}^{t_{m+1}}a(u)[a_{n}((t_{l}-t_{l^{\prime}+1}-t_{m})^{+})-a_{n}((t_{l}-t_{l^{\prime}+1})^{+})]\mathrm{d}u.$
Hence, (74) holds with
$-F_{n}^{(1,3)}(s,r)$ $=\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s\lor r}^{y}[a(u-s)-a(u)][a_{n}((s-r-\Delta_{n})^{+})-a_{n}(\tau_{n}(u)-s)]\mathrm{d}u\mathrm{d}y$
$+\int_{0}^{s}a(u)[a_{n}((s-r-\Delta_{n}-\tau_{n}(u))^{+})-a_{n}((s-r-\Delta_{n})^{+})]\mathrm{d}u\left(1-\frac{s\lor r}{n\Delta_{n}}\right),\,\,$
for $s,r\leq n\Delta_{n},$ and by the arguments used before satisfies that $\rvert F_{n}^{(1,3)}(s,r)\rvert\leq 4a(0)\mathrm{Leb}(A)$ and
$F_{n}^{(1,3)}(\tau_{n}(s),\tau_{n}(r))\rightarrow$ $-\int_{s\lor r}^{\infty}[a(u-s)-a(u)][a((s-r)^{+})-a(u-s)]\mathrm{d}u$
$-\int_{0}^{s}a(u)[a((s-r-u)^{+})-a((s-r)^{+})]\mathrm{d}u,$
as required.
$\ell=1,\ell^{\prime}=4:$
We start by noting that from Lemma 1
$\mathbb{E}(\beta_{j,j,l}^{(1)}\chi_{j}\alpha_{k+1,j}\beta_{k,j,l^{\prime}}^{(1)})$ $=\begin{cases}2\mathbb{E}(\beta_{k,l^{\prime},j}^{(1)2})\mathbb{E}(\alpha_{k+1,j}^{2})&\text{if }k\leq j-l-1;\\ \mathbb{E}(\beta_{k,l^{\prime},j}^{(1)}\beta_{j,l,j}^{(1)})\mathbb{E}(\alpha_{k+1,j}^{2})&\text{if }k\geq j-l;\end{cases}$
$=\mathbb{E}(\beta_{k,l^{\prime},j}^{(1)}\beta_{j,l,j}^{(1)})\mathbb{E}(\alpha_{k+1,j}^{2})(1+\mathbf{1}_{k\leq j-l-1}).$
Thus,
$\sum_{k=l^{\prime}}^{j-1}\mathbb{E}(\beta_{j,j,l}^{(1)}\chi_{j}\alpha_{k+1,j}\beta_{k,j,l^{\prime}}^{(1)})\leq\sum_{i=0}^{j-l^{\prime}-1}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(i,j))\sum_{k=i+l^{\prime}}^{j-1}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k+1,j))\leq(a(0)\Delta_{n})^{2}.$
Therefore, in this case we have that
$\Sigma_{n}^{(1,4)}(l,l^{\prime})=\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}+1}^{n-1}\sum_{k=l^{\prime}}^{j-1}\mathbb{E}(\varsigma_{k,j}\beta_{j,j-1,l}^{(2)})\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))+\mathrm{O}(\Delta_{n}).$
Since,
$\mathbb{E}(\varsigma_{k,j}\beta_{j,l}^{(4)})=\mathrm{Leb}(\cup_{m=j-l}^{j-1}\mathcal{\mathcal{P}}_{A}^{n}(k+1,m))\mathbf{1}_{k\leq j-l-1},$
we conclude that for $j\geq l^{\prime}+l+1$
$\sum_{k=l^{\prime}}^{j-1}\mathbb{E}(\varsigma_{k,j}\beta_{j,j-1,l}^{(2)})\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))=$ $\Delta_{n}\sum_{k=l^{\prime}}^{j-l-1}\int_{t_{j-k-1}}^{t_{j-k}}\left[a(u-t_{l})-a(u)\right]\mathrm{d}ua_{n}(t_{j-k-1}+t_{l^{\prime}+1})$
$=$ $\Delta_{n}\sum_{m=l}^{j-l^{\prime}-1}\int_{t_{m}}^{t_{m+1}}\left[a(u-t_{l})-a(u)\right]\mathrm{d}ua_{n}(t_{m}+t_{l^{\prime}+1}).$
Hence, $\Sigma_{n}^{(1,4)}(l,l^{\prime})=F_{n}^{(1,4)}(t_{l},t_{l^{\prime}})+\mathrm{O}(\Delta_{n}),$ with
$F_{n}^{(1,4)}(s,r)=\frac{1}{n\Delta_{n}}\int_{s}^{n\Delta_{n}-r}\int_{s}^{y}[a(u-s)-a(u)]a_{n}(\tau_{n}(u)+r+\Delta_{n})]\mathrm{d}u\mathrm{d}y.$
Obviously, $\rvert F_{n}^{(1,4)}(s,r)\rvert\leq 2a(0)\mathrm{Leb}(A)$ and by the DCT we also get that
$F_{n}^{(1,4)}(\tau_{n}(s),\tau_{n}(r))\rightarrow$ $-\int_{s}^{\infty}[a(u-s)-a(u)]a(u+r)\mathrm{d}u=:F^{(1,4)}(s,r).$
$\ell=2,\ell^{\prime}=3:$
Plainly
$-\Sigma_{n}^{(2,3)}(l,l^{\prime})=$ $\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}\mathbb{E}(\chi_{k}\beta_{k^{\prime},j-1,l^{\prime}}^{(2)})\mathbb{E}(\alpha_{k^{\prime}+1,j}^{2})\mathbf{1}_{k^{\prime}\leq(k-l-1)\land(k+l^{\prime})}\mathbf{1}_{k\geq l+l^{\prime}+1}\mathbf{1}_{j\geq l^{\prime}+l^{\prime}+2}$
$=$ $\frac{1}{n\Delta_{n}}\sum_{k=l+l^{\prime}+1}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}\mathrm{Leb}(\cup_{i=0}^{k^{\prime}-l^{\prime}}\mathcal{\mathcal{P}}_{A}^{n}(i,k))\sum_{j=k+1}^{n}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k^{\prime}+1,j))$
$=$ $\frac{1}{n}\sum_{k=l+l^{\prime}+1}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}a_{n}(t_{k}-t_{k^{\prime}}+t_{l^{\prime}})\int_{t_{k}-t_{k^{\prime}}}^{t_{k+1}-t_{k^{\prime}}}a(u)\mathrm{d}u+\mathrm{O}(\int_{n\Delta_{n}}^{\infty}a(u)\mathrm{d}u).$
This mean that (74) is fulfilled with
$-F_{n}^{(2,3)}(s,r)$ $=\frac{1}{n\Delta_{n}}\int_{s+r}^{n\Delta_{n}}\int_{s}^{y-r}a(u)a_{n}(\tau_{n}(u)+r)]\mathrm{d}u\mathrm{d}y,$
which is bounded by $a(0)\mathrm{Leb}(A)$ and satisfies that
$F_{n}^{(2,3)}(\tau_{n}(s),\tau_{n}(r))\rightarrow-\int_{s}^{\infty}a(u)a(u+r)\mathrm{d}u=:F^{(2,3)}(s,r).$
$\ell=2,\ell^{\prime}=4:$
Similarly as above we get that
$-\Sigma_{n}^{(2,4)}(l,l^{\prime})=$ $\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}+1}^{n-1}\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}\mathbb{E}(\chi_{k}\varsigma_{k^{\prime},j})\mathbb{E}(\beta_{k,j,l}^{(1)}\beta_{k^{\prime},j,l^{\prime}}^{(1)}).$
In this situation, after some simple change of variables, we infer that
$\sum_{k=l,k^{\prime}=l^{\prime}}^{j-1}\mathbb{E}(\chi_{k}\varsigma_{k^{\prime},j})\mathbb{E}(\beta_{k,j,l}^{(1)}\beta_{k^{\prime},j,l^{\prime}}^{(1)})=$ $\sum_{k=l,k^{\prime}=l^{\prime}}^{j-1}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k^{\prime}+1,k))\mathrm{Leb}(\cup_{i=0}^{(k^{\prime}-l^{\prime})\wedge(k-l)}\mathcal{\mathcal{P}}_{A}^{n}(i,j))\mathbf{1}_{l^{\prime}\leq k^{\prime}\leq k-1}$
$=$ $\sum_{k=l,k^{\prime}=l^{\prime}}^{j-1}\mathrm{Leb}(\cup_{i=0}^{k-l}\mathcal{\mathcal{P}}_{A}^{n}(i,j))\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k^{\prime}+1,k))\mathbf{1}_{k-1\geq k^{\prime}\geq k-(l-l^{\prime})}$
$+\sum_{k=l,k^{\prime}=l^{\prime}}^{j-1}\mathrm{Leb}(\cup_{i=0}^{k^{\prime}-l^{\prime}}\mathcal{\mathcal{P}}_{A}^{n}(i,j))\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k^{\prime}+1,k))\mathbf{1}_{k^{\prime}+1+(l-l^{\prime})^{+}\leq k}$
$=$ $\Delta_{n}[a_{n}(0)-a_{n}((t_{l}-t_{l^{\prime}})^{+})]\int_{t_{l+1}}^{t_{j+1}}a(u)\mathrm{d}u$
$+\Delta_{n}\sum_{m=l\lor l^{\prime}+2}^{j}\int_{t_{m}}^{t_{m+1}}a(u)\mathrm{d}u[a_{n}((t_{l}-t_{l^{\prime}})^{+})-a_{n}(t_{m}-t_{l^{\prime}+1})].$
Therefore, in situation we may choose
$-F^{(2,4)}(s,r)=$ $[a_{n}(0)-a_{n}((s-r)^{+})]\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s}^{y}a(u)\mathrm{d}u\mathrm{d}y$
$+\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s\lor r}^{y}a(u)[a_{n}((s-r)^{+})-a_{n}(\tau_{n}(u)-r-\Delta_{n})]\mathrm{d}u\mathrm{d}y,$
and
$F^{(2,4)}(s,r)=[a(0)-a((s-r)^{+})]\int_{s}^{\infty}a(u)\mathrm{d}u\mathrm{d}y+\int_{s\lor r}^{\infty}a(u)[a((s-r)^{+})-a(u-r)]\mathrm{d}u\mathrm{d}y.$
$\ell=3,\ell^{\prime}=4:$
Finally, in view that for all $l,l^{\prime}\geq 0$, $k\geq l$ and $k^{\prime}\geq l^{\prime}$
$\mathbb{E}[\beta_{k,j-1,l}^{(2)}\varsigma_{k^{\prime},j}]\mathbb{E}[\alpha_{k+1,j}\beta_{k^{\prime},j,l^{\prime}}^{(1)}]\mathbf{1}_{k+1\leq k^{\prime}-l^{\prime}}=\mathbb{E}[\beta_{k,j-1,l}^{(2)}\varsigma_{k^{\prime},j}]\mathbb{E}[\alpha_{k+1,j}^{2}]\mathbf{1}_{k^{\prime}+l+1\leq k\leq k^{\prime}-l^{\prime}-1}=0,$
we conclude that in this situation $\Sigma_{n}^{(3,4)}(l,l^{\prime})=0$.
We are left to show that $\sum_{\ell,\ell^{\prime}=1}^{4}F_{n}^{(\ell,\ell^{\prime})}(\tau_{n}(s),\tau_{n}(r))\rightarrow\sum_{\ell,\ell^{\prime}=1}^{4}F^{(\ell,\ell^{\prime})}(s,r)=\Sigma_{a}(s,r).$ To see this, first note that due to the fact that $\Sigma_{n}^{(\ell^{\prime},\ell)}(l,l^{\prime})=\Sigma_{n}^{(\ell,\ell^{\prime})}(l^{\prime},l)$ and the preceding computations, we have that
$\sum_{\ell,\ell^{\prime}=1}^{4}F^{(\ell,\ell^{\prime})}(s,r)=\sum_{\ell=1}^{4}F^{(\ell,\ell)}(s,r)+\sum_{\ell=1}^{3}\sum_{\ell^{\prime}=\ell+1}^{4}\widetilde{F^{(\ell,\ell^{\prime})}(s,r)}$.
Simple but tedious computations now gives that
$\sum_{\ell=1}^{4}F^{(\ell,\ell^{\prime})}(s,r)=\mathfrak{K}_{4}a(s\lor r)+2a(0)\int_{\rvert s-r\rvert}^{\infty}a(u)\mathrm{d}u,$
as well as
$\sum_{\ell=1}^{3}\sum_{\ell^{\prime}=\ell+1}^{4}F^{(\ell,\ell^{\prime})}(s,r)$ $=\int_{(s-r)^{+}}^{\infty}a(u+r-s)a(u)\mathrm{d}u+a((s-r)^{+})\int_{0}^{(r-s)^{+}}a(u)\mathrm{d}u$
$-\int_{0}^{\infty}a(u)a((s-r-u)^{+}))\mathrm{d}u+\int_{0}^{\infty}a(u+r)a(\rvert s-u\rvert)\mathrm{sgn}(s-u)\mathrm{d}u$
The sought-after identity follows from this and simple algebraic manipulations.∎